### v1

# PHYSICISTS JOKE

Abstract
This work presents a reconstructed English edition of the seminal humor anthology Physicists Joke (Fiziki Shutyat), originally published in the USSR in 1966. While the Soviet edition served as a significant cultural artifact of the 1960s, a large portion of its content consisted of Russian translations of articles originally written in English. To avoid the loss of nuance inherent in back-translation, this edition does not simply translate the Russian text. Instead, the original English source materials have been identified and restored from archives of journals such as Physics Today, The Journal of Irreproducible Results, New Scientist, and Proceedings of the IEEE. The specific structural framework of the 1966 book has been preserved, along with English translations of the unique Russian scientific folklore, anecdotes, and editorial interludes compiled by the original Soviet editors. This collection offers a glimpse into the sociology of science and the shared professional culture of the "Golden Age" of mid-20th-century physics.
To the Reader
The Black Queen* shook her head:
"You may call it ’nonsense’ if you like, but I’ve heard nonsense, compared with which that would be as sensible as a dictionary!"
Lewis Carroll, "Through the Looking-Glass"
*Note: In the Russian translation of Alice, the Red Queen is often called the Black Queen.
The Editorial Board of Physics Literature: It feels somewhat unusual: a humorous book bearing the mark of a scientific publishing house! But this is not accidental. Everything funny in this book will be fully understood and appreciated by those who read serious scientific literature and listen to (and deliver) learned reports—after all, the authors of these jokes are also scientists, sometimes very famous ones.
From this point of view, this book is also scientific. However, it is accessible not only to scientists. Anyone who loves a joke will find pleasure in reading the small humorous pieces collected here, and behind them, they will see serious science—physics—from a new, perhaps unexpected, side. People of all professions love to joke, but these jokes usually do not make it into print and vanish without a trace. It is a pity! Physical folklore is no less interesting than any other; it reflects the history of science, and the life and daily routine of its creators.
What is collected here has been gathered drop by drop. The editors failed to find many things that are known only through oral retelling. But this is a first attempt. There have been no books of this kind until now. Perhaps someday an anthology of physical witticisms (and not only translated ones) will be published; perhaps someone will even write a dissertation on this topic (what don’t they write about!), but for now, the editors have dug through mountains of journals, both printed and handwritten, and gathered the first harvest. As is usually said in prefaces, "let us hope that the collection will please our readers."
Professor Ya. A. Smorodinsky, June 1966
Note on the 2025 English Translation
This book, originally published in the USSR in 1966, was largely a collection of humorous scientific articles translated from English into Russian. To ensure the highest fidelity in this English edition, I have tried not to simply translate the Russian text back into English, which would result in a "broken telephone" loss of nuance.
Instead, wherever possible, I have located the original English source texts (from journals such as Physics Today, The Journal of Irreproducible Results, New Scientist, etc.) and restored the authentic wording used by the original authors. For the short anecdotes, folklore, and specific parodies where the original English text was unavailable or where the Russian version differed significantly from the source, a direct translation from the Russian text has been performed, striving to preserve the humor and style of the 1966 edition.
I would like to acknowledge the assistance of the AI model Google Gemini 3 Pro in locating the original English source texts and translating the specific Russian scientific folklore contained in this volume.
x-Professor Igor Halperin, December 2025
Preface From the Editors
(In Place of a Preface)
— Hello?
— Hello! This is one of the editors of the collection "Physicists Joke." We were recommended to…
— Excuse me, which collection?
— "Physicists Joke."
— What do physicists do?!
— They joke!
— I don’t understand.
— Well, they joke, they laugh.
— Ah, they laugh… Well, so what?
— This will be a collection of translations. Have you or your colleagues happened to come across anything in foreign physics literature…
— No, no! Our staff is engaged in serious business and they have no time for jokes.
…Before we are accused of slandering physicists, let us hasten to assure readers that this conversation was the only one of its kind. Usually—and we sometimes appealed to very busy people—our undertaking was met with full approval and a readiness to help. Physicists value a joke. We believe that in the popular and exciting game of "Physicists and Lyricists," this fact will be counted in our favor.
The idea to compile a real collection matured a long time ago. Reading foreign scientific publications (quite serious ones!), we often found grains, and sometimes whole nuggets of humor which, unfortunately, do not make it into abstract journals or reviews: humorous verses, notes, messages, and even large quasi-serious articles written by physicists and intended mainly for physicists. The question was settled when the volume The Journal of Jocular Physics, published in Copenhagen for Niels Bohr’s 70th birthday, fell into our hands—entirely humorous, something like a printed "skit" written by physicists—Bohr’s friends and colleagues. Having concluded a contract with the "Mir" publishing house, we felt obligated not just to translate the material at hand, but to try to collect the most interesting samples of this peculiar genre.
How to accomplish this task? Bibliographic searches—researching the section "Recreational Physics" in the catalogs of major libraries—turned out to be completely fruitless. "Physics at the Tea Table," "Physics without Instruments," "Physics without Mathematics," and even "Recreational Physics in War"—all this existed, but it was not it. There was no section for "Physics and Humor," and we could only console ourselves with the hope that it might be introduced in the near future…
We had to look through all "suspicious" journals one by one and appeal to colleagues—both acquaintances and strangers. Gradually, material accumulated, very different in character and quality. But still, it turned out to be less than we wanted. And since the criteria we used to decide "include or exclude" were undoubtedly subjective, we remain concerned that the fresh reader’s reaction will resemble the words said by a cafeteria patron to a waiter: "First of all, this is… inedible, and secondly, why is the portion so small?"
All the time we worked on the collection, we were tormented by two problems, which we cowardly put off until the very end. The first problem was the Title. It had to be:
  • Original enough that no one could call it banal;
  • Banal enough that no one could call it pretentious;
  • Pleasing to all editors and translators.
Fortunately for us, this problem resolved itself in the end. It turned out that during the work on the book, one can increase or decrease its volume, change the content… but one thing cannot be done—change its provisional title… for having appeared in advertising brochures, it acquired the force of law.
The second problem was the Preface. Usually, its main content justifies the necessity of publishing the book. But we knew that the scientific necessity of publishing our collection is debatable. And yet we had to justify ourselves somehow: a) to the readers, b) to the publishers, c) to ourselves.
This problem was solved gradually. Point "c" fell away when we decided to compile the collection. Point "b" fell away when the publishing house signed the contract. Point "a" remained, and it caused us the most trouble. It is sad if, when telling a joke, one has to explain the punchline, but it is utterly dreary to explain the punchline before the joke is told. In the end, we decided not to justify ourselves to the readers…
But on one point, we must justify ourselves, or rather, apologize. We offer apologies to the authors of the humorous pieces included in the collection for the inevitable liberties we took during translation. First of all, this concerns cuts in the text. Articles written generally on a serious topic, where the funny parts were just inserts, were shortened significantly. We also excluded places entirely built on untranslatable puns…
Obninsk, May 1965
Yu. Konobeev, V. Pavlinchuk, N. Rabotnov, V. Turchin
Almost Seriously
Dreaming of Physics
American physicist of German origin James Franck (Nobel Prize 1925) once told this story:
"I dreamt the other day about the late Carl Runge, and I asked him: ’How are things in the afterlife? You probably know all the physical laws there?’
He replied: ’Here, we are given a choice: you can either know everything, or you can have the same choice as on Earth. I chose the latter, otherwise it would be too boring.’"
1.1 Physics as a Science and an Art
Author: Karl K. Darrow
Original Source: Physics Today, Vol. 4, No. 11 (November 1951).
The charge of speaking after five such orators as have preceded me is not a light one, and yet is an assignment which should be treated lightly. The hands of the clock are joyously advancing toward the cocktail hour, and they advise me to pervert the famous words beneath a clock in San Francisco and say to myself, "Son, observe the time and fly from wisdom." The organizers of this meeting actually proposed that I should speak under the title "The Whole of Physics". Apart of course from my predecessors on this platform, the last man who could probably have done this was Hermann von Helmholtz. It interests me to realize that there are people still living who studied under Helmholtz; they are the last of our contacts with the era of omniscience. The wishes of the organizers will be formally fulfilled if I succeed in saying nothing that is more irrelevant to any one field of physics than to any other. This condition I will attempt to meet.
I ought to begin with a definition of physics. The American Institute of Physics has provided one, and it would be unseemly to use another in this place. Actually it is a definition of a physicist, but we can easily translate it into a definition of physics. Hearken to it. "A physicist is one whose training and experience lie in the study and applications of the interactions between matter and energy in the fields of mechanics, acoustics, optics, heat, electricity, magnetism, radiation, atomic structure, and nuclear phenomena."
Clearly this is addressed to people who have a clear-cut notion of energy, and therefore not to the general public. But even with respect to its intended audience it has a certain rashness. People who have a clear-cut notion of energy are likely to remember the equation $E=mc^{2}$. This equation operates like a nuclear bomb on the definition, for the definition implies that matter is cleanly and neatly distinguishable from energy, and the equation says it is not so at all. The equation in fact invites us to alter the wording, and say that the physicist is one who concerns himself with the interactions between energy and energy. This has a silly sound, but it is not a silly thought, and I can clothe it in appropriately formal garb by saying that the physicist concerns himself with the interactions between various types of energy. But I will not tamper further with the head of the definition, for it is just the sauce, and the meat is in the tail. Oddly enough, the meat is disguised as a limitation.
There are two limitations here, and one of them is not in Nature and I think that it was not in the minds of the definers. It is implied that in respect of magnetism, for instance, there is one part of magnetism that involves interactions between matter and energy and another that does not. The first part is physics and the second part is not. But there is no second part, and the whole affair reduces itself to a plain and simple definition by enumeration. Physics is a grouping of nine fields like the nine Muses, and the names of the Muses are mechanics, acoustics, optics, heat, electricity, magnetism, radiation, atomic structure and nuclear phenomena. This is what the definers really say and this is the meat of the definition, and all the rest is a valiant attempt to express in a very few words something that slowly dawns on the physicist as he progresses in his science. When in this manner we get down to brass tacks, the only people who can rightly complain are those who would like to have their tacks removed from the list and transferred to some other science than physics, and those others who do not find their tacks in the list and yet would like to be considered physicists. I will not be their spokesman; let them enter their own objections.
The definition also speaks of "study and applications". This sounds like the classic antithesis between pure and applied physics. Let us examine into this distinction, which as will soon appear I deem a necessary evil.
As our science expands, its journals become so huge that they are insupportable in all senses of the word, and the meetings of its cultivators so congested that they defeat their purpose. These are only symptoms: the malady is the finiteness of the human brain, which can absorb only a finite amount of knowledge before old age sets in. But although the malady is incurable the symptoms can be controlled, and this is done by the same technique as prevailed in the cities of ancient Greece and prevails to this day in the beehive. Some of the bees get tired of the congestion and swarm off to another hive. This is the reason and the only reason why the American Physical Society cannot deprecate the newer hives of the Optical Society of America and the Acoustical Society of America, each of which has taken a large piece of physics unto itself. The engineering societies swarmed away a long time ago and they have even larger segments of our domain, but we could not force them back into our hive if we would and we would not if we could. The distinctions are evil in principle, but we cannot get along without them.
Let us try to contrive a definition. One begins by saying that a pure physicist is interested in a device because it illustrates the laws of physics, an applied physicist is interested in the laws of physics because they explain a device. The teacher of physics teaches the dynamo because it exemplifies Faraday’s laws, the teacher of engineering teaches Faraday’s laws because they show how the dynamo works. This definition implies a static science and a static technology. We try to put evolution into it. A pure physicist is one who discovers new laws of Nature, an applied physicist is one who improves an old device or invents a new one.
But many experimental physicists of uncontested purity spend a large part of their time in improving their devices. We must introduce more motive into the definition. A pure physicist is one who improves his devices for no other purpose than to extend his understanding of Nature, an applied physicist is one who improves his devices for any other purpose than to extend his understanding of Nature. On this basis Rutherford was an applied physicist at the start of his career when he was trying to make a radio, purified himself when he abandoned the attempt; Lawrence was a pure physicist until his cyclotrons started to make isotopes which are useful to medical men, then he lost his caste. It is evident that our definition is one of extremes, and it takes a rather single-minded person to hold a position at either extreme. Let us see whether we can discover any analogies in the practice of the arts.
A composer who produces a symphony is presumably a pure musician, one who writes for a dance-orchestra is presumably applied. Yet any conductor knows that the subscribers will not object and will in fact be very pleased if he plays some of the works of Johann Strauss and Manuel de Falla. We are meeting in an opera house. Richard Wagner himself said that the only purpose of his music was to enhance his libretto; he is accordingly an applied musician. Even more singular is the case of Tschaikowsky, who remained a pure musician until he had been in his grave for fifty-odd years, whereupon the sonorous opening theme of his piano concerto in B flat minor was converted into a dance entitled "This Night We Love". I shall leave to people more expert than myself the question whether in the Gilbert-Sullivan team Sullivan was an applied musician or Gilbert an applied poet.
Take painting and sculpture. The pure painter, let us say, is the one whose paintings hang in a museum; the applied painter is the one whose paintings are fitted into the decorative scheme of a house. On this basis Monet and Renoir are applied painters for those who can afford to pay twenty thousand dollars for a picture, pure painters for the rest of us. I do not know quite where to put the portrait painter, except that he is probably pure when his work is hung in a museum with a label "Portrait of a Man". I am reasonably sure that there are many modern painters who, in the inconceivable event that they were present, would wish me to say that the pure painter is the one whose pictures look like nothing on earth, and all the others are applied. There is an analogy to physics in this; we will take another glance at it later.
Architecture ought to be the perfect example of an applied art. Yet I note that there is a doctrine called "functionalism", the exponents of which profess that every part of a building ought to be requisite for its purpose and essential to its structure. The existence of such a doctrine implies that there are buildings with details that are not required by their purpose or their structure, and indeed this is obvious to anybody who has seen a cornice. A drawback of this doctrine is that it forbids you to enjoy a cornice, and indeed in principle it forbids you to enjoy a Gothic cathedral until a civil engineer has proved to you by calculation that if any flying buttress, any pinnacle or any crocket were removed the building would fall down. Then there arises the question of the stained-glass windows: these are functional if they stir a mystical emotion, decorative if they please the tourist, anti-functional if they just impair the light. The first of these views was that of the artists who created the windows of Chartres, the second is that of the guides, the third was that of the eighteenth-century people who improved the lighting by smashing some of the windows and throwing the precious fragments onto the rubbish-heap. It is not easy after all to distinguish what is functional and what is decorative in the totality of a cathedral. A cathedral is a texture of purposeful construction, purposeful decoration, decoration for the sake of decoration, and symbolic instruction. So also is a science. And if some of the sublimest features of a Gothic church derive from the fact that the builders did not have steel beams available, and if the modern builders with steel beams produce a structure that in spite of all its competence is mysteriously lacking in something that we like, these are perhaps analogies with the classical physics and the theories of today.
The Poet
David Hilbert was asked about one of his former students.
"Ah, that one?" Hilbert recalled. "He became a poet. For mathematics, he had too little imagination."
The 440 Volt Student
An interesting example of how words can be used for the quantitative description of measurement results was told by Professor Gale of the University of Chicago.
Professor Gale was working in the laboratory with one of his students, and they did not know what voltage—110 or 220 volts—was on the terminals to which they had to connect their equipment. The student was about to run for a voltmeter, but the professor advised him to determine the voltage by touch.
"But it will just shock me, that’s all!" the student objected.
"Yes, but if it is 110 volts, you will jump back and simply exclaim ’Oh, damn!’, but if it is 220, the expression will be much stronger."
When I recently told this story to students, one of them remarked: "This morning I met a little guy, so he, probably, just before this had connected himself to a voltage of 440!"
1.2 Confessions of an Acoustical Engineer
Author: Marvin Camras
Original Source: IRE Transactions on Audio, Vol. 9, No. 6 (1961).
When I was an assistant, I worked like a horse, and earned so little that I barely had enough for food. My boss always "quipped" about my mental abilities and spent half the day explaining to me what was perfectly clear without him, and then was surprised and threw up his hands that the work was not yet done. He assigned me to develop drawings of unfeasible designs invented by thinkers in our laboratory who were hovering in the clouds. I had to finish and refine everything for everyone to force these constructions to work at least somehow. When I came to the boss with some clever solution, he put it aside and said that this was not what they wanted… Sometimes I worked on a similar project for a year and came to him literally with a masterpiece. Then he would declare: "Very good, my boy, but management decided to tackle a somewhat different topic." Consequently, back to the drawing board…
By that time, as I became a group leader, the situation in the institute had changed. Assistants completely let themselves go and did not know how to do anything, but were offended by every remark. Suppose a trifling job needed to be done. I would have done it myself in a couple of hours. But I had to spend half the day explaining to my assistants why this work should be done at all, why it should be done faster than any other, why it should be done this way and not otherwise, and why necessarily by a certain deadline. When these deadlines passed, I had to listen to half a day of incoherent explanations why the work was not yet done* why nothing works (and, perhaps, will not work) and why we should start everything from the beginning and do everything "as needed". But this "needed" path turned out to be so confused that it usually took at least a year to figure it out, and even less than a year to get some results. At this stage of work somehow, with the help of diplomacy and flattery, persuasion and requests, it was possible to force assistants to make ends meet and push this work off their necks.
Even worse was the case with my administrative duties. I was forced to write so many reports and proposals that this alone would take all my working time. But I still had to answer all current correspondence, answer all phone calls, and receive visitors who came to me in a continuous stream. I had to hire new people and at the same time take care that the old ones remained satisfied with it. It was also assumed that I must take care of the representativeness of the company and for this purpose attend professional meetings, give work in committees, and arrange seminars.
In the end, I started a small, but my own business. Unfortunately, I was unlucky with technical managers and administrators. They have changed much for the worse. They do not strive at all to ensure that something is done by their subordinates. They organize the matter in such a way that everything in the world collapses, and then reorganization begins. They travel, converse, arrange seminars, attend meetings, participate in technical commissions, that is, do anything, but only do not work for the company. The intellectual power enclosed in them and their assistants is colossal, but it is spent not on creation, but on destruction. If the Nobel Prize were awarded for excuses, then our laboratory would have received it long ago.
Now, having gained bitter experience, I dream of becoming an assistant again. An assistant still lives easier. But, unfortunately, I am already married and cannot afford this luxury.
Rutherford’s Method
One evening, Ernest Rutherford entered the laboratory. Although it was late, one of his many students was bent over instruments.
"What are you doing so late?" asked Rutherford.
"I am working," followed the answer.
"And what do you do during the day?"
"I work, of course," the student replied.
"And do you work early in the morning too?"
"Yes, Professor, I work in the morning too," the student confirmed, expecting praise.
Rutherford frowned and asked irritably:
"Listen, when do you think?"
1.3 A Contribution to the Mathematical Theory of Big Game Hunting
Author: H. Pétard
Original Source: The American Mathematical Monthly, Vol. 45, No. 7 (1938).
  1. 1.
    The Inversive Geometry Method. We place a spherical cage in the desert, enter it, and lock it. We perform an inversion with respect to the cage. The lion is then in the cage, and we are outside.
  2. 2.
    The Method of Projective Geometry. Without loss of generality, we can regard the Sahara Desert as a plane. Project the plane into a line, and then project the line into a point inside the cage. The lion is projected into the same point.
  3. 3.
    The Bolzano-Weierstrass Method. Bisect the desert by a line running N-S. The lion is either in the East or West section. Bisect that section E-W. The lion is in one of the two parts. Continue this process indefinitely…
1.4 On the Quantum Theory of the Temperature of Absolute Zero
Authors: G. Beck, H. Bethe, W. Riezler
Original Source: Die Naturwissenschaften, Vol. 19, No. 2 (1931).
We consider a hexagonal crystal lattice. The absolute zero of this lattice is characterized by the fact that all degrees of freedom of the system freeze, i.e. all internal movements of the lattice cease. Naturally, the motion of an electron in its Bohr orbit is excepted. However, according to Eddington, every electron possesses $1/\alpha$ degrees of freedom, where $\alpha$ is the Sommerfeld fine-structure constant. Besides electrons, our crystal contains only protons, for which the number of degrees of freedom is evidently the same, since according to Dirac, a proton can be considered a hole in the electron gas.
To reach absolute zero, we must therefore remove $2/\alpha-1$ degrees of freedom per neutron (which is composed of a proton and an electron). Since $2/\alpha-1$ degrees of freedom correspond to the temperature interval from zero degrees to absolute zero, we obtain for the absolute zero temperature:
$T_{0}=-\left(\frac{2}{\alpha}-1\right)\text{ degrees}.$ (1.1)
Substituting the experimental value $\alpha=1/137$, we obtain:
$T_{0}=-273\text{ degrees}.$ (1.2)
1.5 Earth as a Man-Controlled Spaceship
Author: D. Froman
Original Source: Physics Today, Vol. 15, No. 7 (July 1962).
(Speech at a banquet held after a conference on plasma physics, organized by the American Physical Society in November 1961 in Colorado Springs.)
Since I do not understand plasma physics or thermonuclear fusion very well, I will speak not about these phenomena themselves, but about one of their practical applications in the near future.
Let us imagine that we have managed to invent a spaceship that moves by ejecting reaction products of D-D and D-T. On such a ship, one could travel into space, catch a few asteroids there, and tow them to Earth. (The idea, frankly, is not new.) If we do not overload the rocket too much, it would be possible to deliver 1000 tons of asteroids to Earth, expending only about a ton of deuterium. I, frankly speaking, do not know what asteroids consist of. However, it may well turn out that they are half composed of nickel. It is known that 1 pound of nickel costs 50 cents, and 1 pound of deuterium – about 100 dollars. Thus, for 1 million dollars we could buy 5 tons of deuterium and, having expended them, deliver to Earth 2500 tons of nickel worth 2.5 million dollars. Not bad, right? I was already thinking, should I not organize the American Company for Extraction and Delivery of Asteroids (ACEDA)? The equipment of this company will be extremely simple. With a sufficient subsidy from Uncle Sam, this could be a very profitable business. If anyone present with a large account in the bank wishes to enter the number of founders, let him approach me after the banquet.
And now let’s look into a more distant future. Personally, I generally cannot understand why astronauts dream of getting into interstellar space. In a rocket, it will be terribly cramped. And with food, they will have to cut themselves down significantly. But this is still half the trouble. The main nuisance is that the astronaut in the rocket will be in the same position as a man placed against a beam of fast protons coming out of a powerful accelerator (see figure). I feel very sorry for the poor astronaut; about his sad fate, I even composed a ballad*:
Ballad of an Astronaut
From the beta-inverter and gamma-converter
Remained only the plating.
And the ion cannon, like an empty party popper,
Sticks out, good for nothing.
All mesons decayed, all neutrons decayed,
All visible light radiated away.
According to Coulomb’s law, protons scattered,
For leptons, there is no hope.
The damaged reactor rattles like a tractor,
In the biochamber – rot and mold.
The nozzle is clogged, the bottom is leaking,
And vacuum lashes into the gap…
He flew to Orion, but a flux of gravitons
Crossed his path unexpectedly.
Deviating from course and exhausting all resources,
He barely managed to slip away.
Made a huge hook, flew around half the Universe
And now on an empty ship
In a straight line he was returning home,
Approaching the planet Earth.
But fighting with gravity, super-super-super-acceleration,
He slowed down the hands of the clock.
And the hands froze, on Earth passed
Thousands of thousands of centuries.
Here are the native planets… God! Is this the Sun?
Dark red, slightly warm ball…
Above the Earth it smokes,
Above the Earth swirls Hydrogen, cold steam.
What is this?
Where is the human tribe? – In unknown, distant worlds.
Their children grow
Already on a new planet, And Earth is all in cosmic ice.
Swearing and crying
From such a failure, the Astronaut turned the lever,
And B rang out
And A rang out
And X rang out –
BANG!
* Free translation from English by V. Turchin.
But I feel sorry for those who will remain on Earth. After all, our Sun is not eternal. It will someday fade, plunging everything surrounding it into cosmic darkness and cold. As Fred (Fred Hoyle, that is)* told me, in a couple of billion years on Earth it will be so cold that not only comfort, but life itself on this planet will be out of the question. And consequently, it makes explicit sense to move somewhere. It seems to me that for the majority of us, the most convenient spaceship would be the Earth itself. Therefore, if we do not like that our luminary is gradually fading, and generally if we get tired of everything in the Solar System, why remain here? Let’s fly somewhere directly on our Earth. In this case, all difficulties associated with space flight will fall away by themselves. After all, the problem of protection from radiation does not exist, there is an atmosphere on Earth, and the speed of movement will be small. Safety and the pleasantness of such a journey are obvious. However, will we have enough energy?
* Fred Hoyle – famous English astrophysicist, professor at Cambridge University, author of a number of works on theoretical astrophysics, cosmogony, theoretical gravity and… several science fiction novels.
First of all, heat and light will be needed: after all, during the long time we will be removed from the Sun or any other star. Deuterium contained in ocean water can give us $10^{38}$ ergs, consequently, if we use it only for heating and lighting, then this will be enough for three million years – a term quite sufficient. True, here is a small snag. At our speeds, we will consume $3\dots 10^{10}$ pounds of deuterium per year, and the cost of it is 100 dollars per pound, consequently, the consumed deuterium will be 100 times the annual budget of modern air forces. But, perhaps, it will be possible to obtain deuterium at wholesale prices?
However, we will need more energy to tear away from the Sun. Calculation shows that for this, $2.4\dots 10^{40}$ ergs will be needed, that is, much more than all the ocean deuterium can give. Therefore, it will be necessary to find other sources of energy. I suggest that for the solution of this problem we will have to turn to the synthesis of an alpha-particle from four protons. When using this reaction, all protons of the world ocean will give us energy $10^{42}$ ergs, that is, 40 times more than is needed to tear away from the Sun.
As a working body, sand can be used. Ejecting 1000 molecules of $SiO_{2}$ for every synthesized alpha-particle, we will need to spend only 4% of the Earth’s mass to tear away from the Sun. It seems to me that we can afford this. Especially since for such a purpose, it is not a pity to use the Moon: after all, far from the Sun, it makes no difference to us anyway. Leaving the Solar System and wandering in cosmic space, we, probably, will be able to replenish our reserves of mass and energy from time to time, refueling on the fly at the expense of planets encountered on the way. On the path of carrying out these plans, there is currently one fundamental obstacle: we do not know how to carry out the chain reaction $4p\to He^{4}$. Now you see what an important problem this is. We need to double our efforts to solve it. Time does not wait: Earth has spent already two-thirds of the term allotted to it by the Sun.
I assure you: in space, we will be excellent. Perhaps, we will like it so much that we won’t even want to attach ourselves to a new star.
Einstein and the Hat
Einstein was visiting acquaintances. It started to rain. When Einstein was about to leave, they offered him a hat.
"Why?" said Einstein. "I knew it would rain, and that is exactly why I didn’t wear a hat. It dries longer than my hair. That is obvious."
The Selection Process
"I simply cannot find myself an assistant," Thomas Edison once complained to Albert Einstein. "Young people come in every day, but not a single one is suitable."
"And how do you determine their suitability?" Einstein inquired.
Edison showed him a sheet of paper with a list of questions. "He who answers these will become my assistant."
Einstein picked up the list. "’How many miles is it from New York to Chicago?’" he read, and immediately answered: "One needs to look that up in a railroad guide."
"’What is stainless steel made of?’" Einstein continued reading. "One can find that out in a metallurgy handbook…"
Running his eyes quickly over the remaining questions, Einstein handed the paper back and said: "Without waiting for a refusal, I withdraw my candidacy myself."
1.6 List of Typical Examination Questions for Physics Graduate Students
Author: Harry J. Lipkin
Original Source: The Journal of Irreproducible Results, Vol. 7, No. 2 (1959).
  1. 1.
    MECHANICS. A particle moves in a potential field $V(r)=e^{-r/r^{12}}$.
    1. (a)
      Show that the solution to this problem bears no relation whatsoever to the binding energy of the deuteron.
    2. (b)
      Explain the asymptotic behavior of the solution as $r\to\infty$.
  2. 2.
    ELEMENTARY PARTICLES. List all elementary particles that have not yet been discovered, indicating their mass, charge, spin, isotopic spin, strangeness, and the reasons why they have not yet been discovered.
  3. 3.
    QUANTUM THEORY. Write down the Schrödinger equation describing a student studying the physics of elementary particles. Derive an expression for the operator "Passed – Failed", which has the eigenvalue +1+1 if the student passes the session, and 1-1 if they fail. Show that the state of the student at the end of the semester is always an eigenstate of this operator.
  4. 4.
    SYMMETRY PROPERTIES. Investigate the properties of the Dirac equation with respect to rotation:
    1. (a)
      When the blackboard on which the equation is written is rotated;
    2. (b)
      When the physicist investigating the equation is rotated.
  5. 5.
    NUCLEAR REACTIONS. A coin enters into interaction with a machine selling Coca-Cola. Determine the relative probabilities of the following reactions:
    1. (a)
      Capture (input channel: nickel*, output: nothing);
    2. (b)
      Elastic scattering (n, n) (input channel: nickel, output: nickel);
    3. (c)
      Reaction (n, 2n) (input channel: nickel, output: two nickels);
    4. (d)
      Reaction (n, p) (input channel: nickel, output: button);
    5. (e)
      Reaction (n, c) (input: nickel, output: Coca-Cola).
    *Note: Nickel = 5 cent coin.
  6. 6.
    RELATIVISTIC QUANTUM FIELD THEORY. A pair of twins, Bingle and Dingle, is born. Immediately after birth, Dingle is sent in a rocket toward a distant star at a speed of $0.999c$, while Bingle remains to work in the laboratory. Determine the relative age of Bingle and Dingle at the moment of Dingle’s return, taking into account the following process: at the farthest point of his trajectory, Dingle emits a virtual pi-meson, which creates a Bingle-Anti-Bingle pair. The Anti-Bingle returns to Earth, where it annihilates with Bingle, while Dingle and the new Bingle happily live out their days near the distant star.
  7. 7.
    EXPERIMENTAL TECHNIQUE. Describe the most expensive method for determining Planck’s constant.
  8. 8.
    DISPERSION RELATIONS. Explain the phenomenon of the multiple production of strange articles in nuclear physics, which are observed in the non-physical region of Physical Review. Show that the principle of causality allows one to fully predict the results of any experiment, with good agreement observed up until the moment someone actually performs the experiment.
1.7 Typology in Scientific Research
Authors: A. Kohn and M. Breier
Original Source: The Journal of Irreproducible Results, Vol. 7, No. 2 (1959).
  1. 1.
    THE DISCOVERER (The Idea Man). These scientists are always "giving out" new ideas. Their brains are always ready to sink their teeth into random prey. Good scientific training allows them to quickly assess the importance of an observed fact and formulate an idea, after which the hypothesis is ready (or at least a working hypothesis). Then they either verify it experimentally themselves or leave it to others to worry about, deriving pleasure from the purely speculative solution of the problem.
  2. 2.
    THE EXPLOITER. These are researchers with a quick grip; their eyes and ears are constantly open. Such a scientist is rarely found in his own laboratory; he prefers to spend his time in discussions with colleagues from other laboratories and institutes, especially if these colleagues are working on something that interests him. He never lacks good ideas, which, although they were not born in his head, turn into interesting articles, generously sprinkled with references to "private communications."
  3. 3.
    THE CONNOISSEUR (The Critic). The mental abilities of such a man far exceed his capabilities (and desire) to conduct his own experiments. He is capable of evaluating (and does evaluate) good work, and often does this better than the author himself. A critical mind, combined with innate inconsistency, is the reason why the results of each subsequent series of measurements significantly differ from all previous ones; this prevents such a scientist from publishing anything unless he has a decisive boss.
  4. 4.
    THE IMPROVER. He resembles the "Connoisseur," but possesses slightly higher productivity. His achievements are represented by very few, but excellent articles, based on experiments which he repeated so many times that all unexpected or unpredicted results managed to be discarded with the help of sophisticated statistical processing.
  5. 5.
    THE MAN ON THE LEVEL (The Well-Read Man). He knows everything that is worth knowing. Unlike the "Exploiter," he spends all his time in the library, where it is rare for anyone to beat him to receiving the latest issue of a journal.
  6. 6.
    THE CO-AUTHOR. This type has perfectly mastered the art of scientific diplomacy. He achieves the inclusion of his name in the lists of authors of most articles published by employees of the department where he works, without any pressure, and his contribution is sometimes expressed merely in deciding a question—whether to use the conjunction "and" in the title of the article. Some people hold the opinion that the "Co-author" is almost the same as the "Advisor," but what an advisor is, everyone who plays cards or chess knows.
  7. 7.
    THE ADVISOR (The Consultant). Advisors, who are found most often in institutions engaged in fundamental research, should not be confused with consultants who, engaged in science, give advice to the relevant branch of industry, giving people appropriate advice for appropriate remuneration.
  8. 8.
    THE GADGETEER. In the world of modern science, there is research, the conduct of which is absolutely impossible without a solid set of instruments. Opportunities that were available to Archimedes and Newton are now limited to no one (except, perhaps, theoretical physicists), and a well-equipped laboratory is as necessary for productive research as a typewriter is for writing reports to management. Therefore, some scientific workers consider the meaning of their life to be obtaining and (if possible) using the largest possible number of extremely modern instruments. Visiting such a laboratory is a sheer pleasure. The soul simply rejoices at the sight of rooms packed with ultra-modern equipment that sparkles with glass and nickel; awe-inspiring names of numerous installations are used more to impress visitors than for any other purpose.
  9. 9.
    THE PUBLISHER. By this term, for lack of a better one, we will denote those who, by any numerous means at an exponentially increasing rate, lengthen the list of their scientific works. A variation of this type is:
    • The Re-hasher (The Ruminant) (a name associated with the peculiarities of the digestive process in certain mammals). Such a person usually settles in some underdeveloped country. He publishes his observations there and findings, prefacing them with the introduction: "For the first time in history… (insert name of country) it was observed…", after which a translation into the local language of some work done by other people in another place is honestly reproduced.
    Another representative of the same type:
    • The Multiplier. An individual who lays out the results of his work or his speculative reasoning into the largest possible number of small packets with labels "Article," "Letter to the Editor," "Short Communication," etc., and successfully increases the list of his works in this way.
  10. 10.
    THE CORRESPONDENT. To get acquainted with this type of scientist, we refer the reader to the last pages of any scientific journal, where headings such as "Notes," "Brief Communications," "Preliminary Results," or "Letters to the Editor" are dazzling. There, every Correspondent reports something truly important, very similar to a great discovery, which follows to publish as quickly as possible before someone else does… Such messages end with the words: "A detailed description of experiments (or results) will be published in such-and-such an edition (or in the near future)." In 50% of cases, the promised publication never appears, since the results of repeated experiments beat off any interest in the idea itself.
The above list by no means claims to be complete. In the literature (in particular, in the monograph by Birch) one can find excellent descriptions of The Re-discoverer, The Continuer, The Thinker, The Spreader, The Talker, The Pusher, The Imposter, The Disqualifier, and many others. One should also remember the existence of The Destroyers, The Refuters, The Enthusiasts, The Neglecters, The Complicators, etc. We are sure that a reader with imagination will easily be able to construct images of all the scientists with whom he is personally acquainted.
Contributions from "The Journal of Jocular Physics"
(Jubilee Collections in Honor of Niels Bohr)
2.1 A New Tale About the Curious Elephant
(A parody of Rudyard Kipling’s "The Elephant’s Child")
No, this is not the tale about that bad Elephant that Kipling wrote about… this is a tale about another, good Elephant… who lived almost in all countries of the world. This Curious Elephant had a wonderful nose from birth, so he did not need the services of the Old Crocodile (Rutherford), and over time he discovered a new thing—the Atomic Age.
2.2 The Atom That Bohr Built
(A parody of "The House that Jack Built")
This is the Atom that Bohr built.
This is the Proton, Placed in the center of the Atom, That Bohr built.
And here is the Electron, That rushes around the Proton, Placed in the center of the Atom, That Bohr built.
Here is the Mu-Meson, That decayed into an electron, That rushes around the Proton, Placed in the center of the Atom, That Bohr built.
And here is the Pi-Meson, That, decaying, gave a Mu-Meson, That decayed into an electron…
And here is the Bevatron, In which that Proton was accelerated, Which in a collision gave birth to a Pi-Meson, Which, decaying, gave a Mu-Meson…
This is the Law, Proclaimed by Bohr. A Law for all nations, A Law for all times, Successfully describing from two sides Not only the Proton And Electron, But also the Neutron, Photon, Positron, Phonon, Magnon, Exciton, Polaron, Betatron, Synchrotron, Phasotron, Cyclotron, Cyclone, Ceylon, Nylon, Perlon, Cologne, Decameron.
And, undoubtedly, every neuron of the Brain, By which That remarkable Bevatron was invented, In which that Proton was accelerated, Which in a collision gave birth to a Pi-Meson, Which, decaying, gave a Mu-Meson, That decayed into an electron, That rushes around the Proton, Placed in the center of the Atom, That Bohr built!
2.3 On the Standardization of Scientific Papers
Original Source: Confidential CERN/T/000. Distributed to all CERN members.
Title: On the Question of             in the Generalized Nuclear Model.
Authors:                and               .
In view of the serious difficulties arising from the attempt at a precise description of the properties of an ensemble of strongly interacting particles, we consider the following approximate Hamiltonian:
$\underline{\hskip 113.81102pt}\dots\underline{\hskip 113.81102pt}\dots,\quad(1)$ (2.1)
where             denotes the corresponding generalized coordinates. The Hamiltonian, thus, consists of three terms:                describes collective motion,             – the motion of individual particles, and                – the         (strong / weak / intermediate – cross out unnecessary) interaction between them.
For the energy of low-lying excited states, we thus obtain:
$\underline{\hskip 113.81102pt}\dots\underline{\hskip 113.81102pt},\quad(2)$ (2.2)
which corresponds, of course, simply to the product of $\hbar^{2}/2m$ and $i(i+1)$, as was to be expected [1]. The system does not possess central symmetry, which allows us to describe its surface as a deformed sphere. The moment of inertia, consequently, is determined by the pole of the expression               , which leads to the formula:
$\underline{\hskip 113.81102pt},\quad(3)$ (2.3)
where, however, the dependence of the parameter             on             is unknown.
The authors are deeply grateful to Director                for his displayed interest in the work. One of us (           ) is highly thankful to            .
2.4 The Key to the Key System
A letter to the editor.
Earlier, an opinion was expressed that the system of door keys in our institute is more complex than field theory. This is a clear distortion of facts, and to refute it, in this communication we present a simplified theoretical scheme on the basis of which this system was created.
Let us introduce three axioms:
  1. 1.
    Pins turn levers; for the lock to open, all levers in the lock must be turned.
  2. 2.
    If there is no pin, hole, or lever in a given position, we will speak of an anti-pin, anti-hole, or anti-lever.
  3. 3.
    There are no levers behind anti-holes, for such a lock could not be opened.
…Using this formalism, it is easy to find the total number of keys opening a given lock.
The Population of Italy
Found on the wall at the University of Milan Physics Department.
Total Population of Italy 52,000,000 Over 65 years old 11,750,000 Remaining for work 40,250,000 Under 18 years old 14,120,000 Remaining for work 26,130,000 Non-working women 17,315,000 Remaining for work 8,815,000 University students 275,000 Remaining for work 8,540,000 Sick, insane, vagrants 1,310,000 Remaining for work 1,240,000 Illiterate, artists, judges 880,000 Remaining for work 360,000 Ministers, deputies, senators 119,998 REMAINING FOR WORK 2
Who are these two? You and Me.
Let this tragic reality serve as a signal of alarm… especially for
You, because I am tired of doing my duty to the country alone.
Useful Advice
3.1 How to Write Scientific Articles
Author: L. Solymar
Original Source: Proceedings of the IEEE, Vol. 51, No. 4 (April 1963).
(Original Title: "Getting into the Act")
Introduction
The question of preparing scientific papers for publication has been discussed from various points of view, yet many of its aspects still remain neglected. It is also surprising that the great advances achieved in the last decade in the conduct of scientific research have brought us almost no closer to a final solution of this problem. Many books and brochures have been written on "how to write articles," but they are all devoted either to vague general recommendations ("write clearly," "explain your thoughts," "do not deviate from the topic," etc.) or to advice on technical formatting ("leave margins on one side," "captions for figures must be typed," "illustrations should not exceed $10\times 15$ cm," etc.). Without denying the seriousness and importance of this advice, I nevertheless believe that it touches only upon a limited circle of secondary questions. In this note, I do not intend to introduce new ideas, but simply wish to share my experience in compiling technical articles and the valuable comments I have received from friends and acquaintances.
Some Considerations on the Motives Prompting the Writing of Articles
A whole range of reasons (from ordinary graphomania to the desire to improve one’s social status) induces a person to write and publish their scientific works. I will not go into details and will limit myself to considering only four main motives: 1) Selfless desire to spread knowledge; 2) Concern for one’s own priority; 3) Concern for one’s professional reputation; 4) Desire for promotion.
Under the influence of the first reason, chiefly young people write, and even then, apparently, only when preparing their first scientific work. The number of such authors is small, and for the majority of them, the first article is also the last. Consequently, the first reason cannot be placed in the same rank as the others, which are more powerful, although one should not forget it entirely.
The second reason—priority—moves only a small group of authors, although in importance it far surpasses any other reason. The desire to link one’s name with some discovery has long been a distinctive feature of scientific workers. Since publication began to serve as proof of discovery, there has been a desire to forge one’s articles as quickly as possible. However, the author must not forget the possibility of further use of his discovery. If he publishes the data he has obtained, someone might carry these ideas to the end and deprive the author of the opportunity to reap the fruits of his labor. The ideal solution to the question is to guarantee priority by announcing the discovery, but to delay detailed publication until the full assessment of its potential possibilities. As is known, the first scientist to apply this method was Galileo Galilei, who sent a description of his astronomical discoveries to Kepler in the form of an anagram, and deciphered its content only a year later. Since modern scientific journals, unfortunately, do not usually publish anagrams, current discoverers (or inventors) must act in a different way. I would recommend starting articles with an intriguing title, for the greater the impression the title makes, the less information needs to be communicated in the article itself. For example, a headline "Amplifier with Loaded Negative Inductance" will immediately convince every reader that an important new principle has been discovered. The author will be forgiven if he is rather vague on the essential points and reports on the discovery only in general terms.
The third reason is concern for professional reputation. A high professional reputation can be achieved in various ways. It is enough, for example, to make an outstanding invention or, even better, to receive a Nobel Prize, and your competence in this matter will be beyond doubt. However, for the overwhelming majority of scientific workers, the only accessible way is to write a large number of articles, each of which contributes at least a small contribution to science. It is advisable to limit the first few articles to a narrow topic (for example, "Connections in Waveguides") to win recognition. However, later the author must testify to his versatility by writing several works covering a broader topic (for example, "Ultra-High Frequency Oscillations"). After publishing three dozen articles, fame reaches saturation and will no longer grow with a further increase in the number of printed works. Here comes the most suitable moment to stop printing (a few review articles do not count) and attempt to take a leading administrative position.
Advice on Formatting Manuscripts
So far I have considered only the reasons why scientific papers are written. Now I would like to touch upon the position of a young author (who does not have powerful co-authors), whose article is to pass through the gauntlet of reviewers.
How to ensure the acceptance of an article for publication? Usually, reviewers are selected from among leading scientists to filter out from the general flow of manuscripts those that are worth printing (after editing). Unfortunately, leading scientists, as a rule, have little time, and many responsibilities, and in addition, they bear the burden of administrative cares. They cannot devote the bulk of their time to reading some article, and yet they must make critical remarks.
A beginner author should take this circumstance into account and, in order not to waste time later on complaints, he needs to write his article in such a way that it satisfies the requirements of the reviewer, whose sharp eyes will detect the slightest anomaly. If the article is too long, the author will be accused of verbosity; if the article is too short, he will be advised to collect additional material. If he reports on purely experimental work, he will be subjected to criticism for "lack of theoretical justification"; if he offers a theory for discussion, he will be called "superficial." Therefore, I propose a compromise. The article should have a volume of 8 to 12 pages, typed, and about one-third of it should be occupied by mathematical formulas.
Following the above recommendations, the author has a fair chance that his article will pass independently of its content. A quick glance at such an article will evoke the reviewer’s favor. Next, everything depends on his reaction within the next thirty minutes. If during this time he can quickly make critical remarks on three minor errors, the article will be accepted.
Thus, the main task of the author is to give the reviewer material for three minor remarks. Below we provide several recommendations for facilitating the choice of such material.
  1. 1.
    Pick an unsuccessful title (all reviewers love to suggest their own titles).
  2. 2.
    "Forget" to define one of the notations in the very first equation.
  3. 3.
    Make a spelling mistake in a word (only one!), which is often spelled with a mistake.
  4. 4.
    Deviate from standard notations (we are talking only about one parameter).
  5. 5.
    Write exp x and $e^{x}$ interchangeably.
I hope that the remarks I have made will contribute to a better understanding of the essence of the work on compiling a scientific article and at the same time serve as a guide for beginner authors.
3.2 Glossary for Research Reports
Author: C. D. Graham Jr.
Original Source: Metal Progress, Vol. 71, No. 5, pg. 75 (1957).
THEY WRITE THEY MEAN
It has long been known that… I haven’t bothered to look up the original reference.
…of great theoretical and practical importance. …interesting to me.
While it has not been possible to provide definite answers to these questions. The experiments didn’t work out, but I figured I could at least get a publication out of it.
The W-Pb system was chosen as especially suitable to show the predicted behavior… The fellow in the next lab had some already made up.
High purity…
Very high purity…
Extremely high purity…
Super-purity…
Spectroscopically pure…
Composition unknown except for the exaggerated claims of the supplier.
A fiducial reference line… A scratch.
Three of the samples were chosen for detailed study… The results of the others didn’t make sense and were ignored.
…handled with extreme care during the experiments. …not dropped on the floor.
Typical results are shown… The best results are shown…
Although some detail has been lost in reproduction, it is clear from the original micrograph that… It is impossible to tell from the micrograph.
Presumably at longer times… I didn’t take the time to find out.
The agreement with the predicted curve is excellent. Fair.
Good. Poor.
Satisfactory. Doubtful.
Fair. Imaginary.
…as good as could be expected. Non-existent.
These results will be reported at a later date. I might get around to this sometime.
The most reliable values are those of Jones. He was a student of mine.
It is suggested that…
It is believed that…
It may be that…
I think…
It is generally believed that… I have such a good objection to this answer that I shall now raise it.
It is clear that much additional work will be required before a complete understanding… I don’t understand it.
Unfortunately, a quantitative theory to account for these effects has not been formulated. Neither does anybody else.
Correct within an order of magnitude. Wrong.
It is to be hoped that this work will stimulate further work in the field. This paper isn’t very good but neither are any of the others on this miserable subject.
Thanks are due to Joe Glotz for assistance with the experiments and to John Doe for valuable discussions. Glotz did the work and Doe explained what it meant.
3.3 How to Address the American Physical Society
Author: Karl K. Darrow
Original Source: Physics Today, Vol. 14, No. 10 (October 1961).
Consider an actor in a hit show on Broadway, and contrast him with a physicist addressing the American Physical Society. The actor has all the advantages. He is speaking lines written for him by a master of the art of commanding the interest of an audience (remember that we are postulating a hit show). He has a gift for acting, and also a long experience in the art; otherwise he would not be in the cast. Even so, he is not allowed to speak his lines in any way that occurs to him. Every phrase, every inflection, every gesture, even the position that he is to take on the stage, has been tested or even prescribed by a professional director, who does not hesitate to give him mandatory instructions, or even to alter the lines if they seem ineffective.
One might assume that assured of such splendid collaboration, the dramatist would write a play two hours long without a break, and the manager would be content to offer the play in a barn with benches for the seats. This is apparently not the view of those who are experienced in such matters. Ample intermissions are provided, and an act which runs for as much as an hour is sufficiently rare to cause the critics to mention it. Usually the theatre has comfortable chairs and is well ventilated, or even air-conditioned. All this is provided to induce people to come to a play for the apprehension of which, with rare exceptions, no intellectual effort is demanded.
Now consider the physicist. He has thought out his own lines, and is not always proficient in this not altogether easy art. He has little or no training in the art of elocution, and no director has rehearsed him. His subject requires a considerable amount of mental effort on the part of his listeners. The listeners themselves are usually uncomfortable and sometimes acutely so. This may be because the chairs are uncomfortable, or because the room is hot and stuffy, or because the program has already been running for an hour or more without a break; or two or all three of these conditions may exist together. Laurence Olivier or Helen Hayes might well quail at the prospect of having to sway an audience under such conditions. Under these highly unfavorable circumstances, does the physicist strive to put on a reasonable facsimile of Olivier or Hayes? It may be conjectured that frequently he does not, because of the popularity of the saying that when a meeting of the American Physical Society is going on, the members are in the corridors or on the lawn instead of listening to the speakers. People with tickets to Turandot are not standing around on the sidewalks outside of the Metropolitan Opera House when the curtain is up.
Can anything be done to amend this situation? Very little, I am afraid; but the following suggestions point in the right direction.
  1. 1.
    Speak loudly enough to be heard in the remotest part of the room. Some people sincerely believe that their voices are too weak to achieve this. No doubt this is sometimes the case, but I venture to believe that most of them are wrong. In my youth I was constantly reproached for speaking too faintly, and I thought that I could not help it; experience proved me wrong. I do not think that I could manage a speech in the Metropolitan Opera House without an amplifier, but a physicist is not likely to be asked to speak in so large a hall, and if he were he could count on the presence of an amplifier. In a hall seating three hundred persons or fewer, the amplifier ought to be unnecessary except in pathological cases. If there is an amplifier, do not expect it to transform a conversational tone into a loud one. It is better to go to the opposite extreme, and pretend to yourself that the microphone is not there, even though you are speaking directly into it.
    The trick recommended by those who instruct speakers is to look at and speak to the people in the rear row. This is often made difficult by the fact that some of the prominent people in the audience are sitting in the front rows; this is particularly common in University colloquia. If this situation exists, ignore it. If Niels Bohr is sitting in the front row and Joe Doakes in the rear row, speak to Joe Doakes. Bohr will hear you.
  2. 2.
    Write out your speech in advance, and commit it to memory. I have heard only one objection (from the viewpoint of the audience) raised against this procedure, and it seems to me groundless. It has been contended that a written speech is dull and lifeless; the implication is that an unwritten speech glitters with sparkling impromptus. But the presence of a manuscript need not prevent the speaker from substituting a sparkling impromptu for something that he has written; and if the impromptu fails to occur to him, the manuscript is there to carry him along. Of course, it is possible to memorize a speech without writing it out; this is recommended to those who hate to write. It is a fact that a good speech is likely to be looser in texture than a good article. No difficulty will arise from this cause if the speaker remembers that it is a speech that he is writing.
    There are some who think that it is better to hear an unprepared physicist groping for what he wants to say than a prepared physicist saying what he wants to say. It would be fascinating to see this theory given a trial by the Royal Festival Ballet, but nobody ever will. For an advanced student of the dance it may be instructive to see a dancer fall on her face, pick herself up, and resume her part in the ballet; but for practically everyone else it is acutely embarrassing.
  3. 3.
    If you cannot memorize your manuscript, read it aloud. This bit of advice will probably be resented, for we have all suffered from dreary speeches poorly read. There is, however, no compelling reason why a manuscript should be poorly read. Lady Macbeth has to read a letter aloud in an early scene of the play; it is one of the high points of the drama. More than forty years ago Ethel Barrymore read a letter aloud in such a way that it is still remembered by elderly playgoers, though the play itself is forgotten. The trouble is largely that most readers glue their eyes to the manuscript for seven-eighths of the time, lifting their eyes from time to time to steal a glance at the audience as though to make sure that it is still there. Reverse the ratio. It is easy to keep your eyes on the audience during seven-eighths of the time and look at the manuscript during the other eighth. For a manuscript which you have composed yourself, it should be extremely easy. Try it and see.
  4. 4.
    Situate your topic in the general framework of physics at the beginning, and summarize your conclusions at the end. Even in a ten-minute paper, a minute at the beginning and a minute at the end are not too much to reserve for these purposes. Do not fear to repeat your main points. I shall have more to say on this topic of repetition near the end.
  5. 5.
    Time yourself. The members of the American Physical Society are now pretty well trained in the art of giving ten-minute papers, but longer ones are still apt to overrun. This is particularly serious when the closing bell rings when the speaker still has five minutes to go, and these five minutes comprise the conclusions which are the incentive for the paper. The speaker naturally does not want to omit the climax of his speech, and the chairman is seldom ruthless enough to insist.
    This is where a manuscript is particularly useful. Timing-marks can be inserted at the end of each page or along the margin, and the speaker (who should constantly be looking at his watch) will then know when he is running behind and will be able to catch up by leaving out relatively dispensable passages. One hundred and thirty words a minute, or say two-and-a-half minutes for a double-spaced typewritten page, is fast enough. In the timing, allow for twenty seconds or thereabouts of silence just after you make each of your difficult points. These gaps will give the audience a chance to think about what you have said; there are no laws requiring a speaker to be talking all of the time at his disposal. The difficulty in timing is greatest when the paper involves blackboard work or slides. Rehearsal is necessary in such cases, and is worth the effort.
  6. 6.
    Aim your discourse toward the average of the audience, not toward the topmost specialists. Too many young theoretical physicists speak as though they were instructing Oppenheimer; too many band-spectroscopists, as if they were addressing Mulliken; too many solid-state physicists, as though the audience consisted of Seitz—and so it goes. This is not quite so flagrant a fault as it was in the days before the meetings of the Society splintered into simultaneous sessions, each attracting its own coterie of specialists; but it is still an error, and anyone who avoids it is doing his bit toward the all-important end of keeping physics from breaking up into a horde of narrow specialties.
    There is one specious argument for the procedure which I am deprecating here. The young man may think that the topmost specialist is also the prime job-giver, and therefore is the man whom it is urgent to impress. But in the first place, it seems plausible to suppose that the topmost specialist forms his opinions of the neophytes from their writings and from personal contacts; and in the second place, the job-giver in the audience may be, say, some chairman of a department of physics whose own specialty lies elsewhere, and who is going to assess the young man by his lucidity and not by his profundity. If these entirely reasonable suppositions are correct, the young man is doing himself a disservice by speaking as though he were addressing exclusively those who know more than he.
  7. 7.
    The problem of the blackboard. This is one of the toughest of all problems, and here the theatre is of no use. I have never seen a play in which an actor had to write on a blackboard. I think that an actor would write on the blackboard without saying a word, and then turn to the audience and speak. For a physicist the psychological inhibition against doing this is quite invincible, but at least the attempt should occasionally be made. He can at least avoid the tendency to drop the level of the voice while addressing the blackboard. There are, however, two faults at the blackboard which can often be avoided.
    One should write his symbols large enough so that they can be read from the back of the room. I hope I never forget the shock which I once experienced when, having finished what I had fondly supposed to be a good lecture, I went to the back of the room and found that nothing I had written could be read beyond the middle rows. Sometimes the speaker finds the blackboards to be much smaller than he had reasonably counted on; in such a case he has to choose between altering his presentation and confining his effectiveness to the people in the nearer rows. Sometimes, of course, either the chalk or the blackboard is impossibly bad; the speaker is then helpless unless he is good enough to revise his plans and do the whole speech without the blackboard. One ought also to write his equations in the order in which he speaks them, instead of putting each in the nearest convenient empty spot and dabbing with the eraser to make more empty spots, so that at the end the board is littered with incoherent symbols. One should know in advance just how the board will look at every moment during the discourse, and at the end of the talk the board should carry all of the principal equations arranged in logical order. I am afraid that this is a counsel of perfection.
  8. 8.
    The problem of slides. Most people who show slides at all show too many and show them too fast. (I suspect that this is often because the speaker has prepared too long a speech and tries to compensate by racing through the slides.) Rare is the slide which can be properly apprehended in less than thirty seconds, though exceptions do occur. It is impossible to assign a rigid maximum to the number of slides which can be shown effectively. I suggest seven for a ten-minute paper, but I make exception for the cases in which the argument is shown on slides instead of on the blackboard. The one advantage of the blackboard over slides is that the overfast speaker is obliged to slow down as he writes; this advantage can be shared by the slides if the speaker will give them time enough. There is much else excellent advice to be given about slides, but it has all been said by J. R. Van Pelt in the July 1950 issue of the American Scientist. This should be required reading for all physicists.
  9. 9.
    The problem of the "jargon". Some people ascribe the difficulty of understanding science to what they call the "jargon". This seems to imply that scientists use long technical terms out of perversity, when they could just as well use short familiar words. This is absurd. If I am giving a speech on a subject involving entropy or a synchrocyclotron, less than nothing will be gained if I avoid the word entropy or the word synchrocyclotron by some cumbrous periphrase or by some vivacious popular word which does not mean the same thing. Entropy is entropy and a synchrocyclotron is a synchrocyclotron, and there is no synonym for either. On the other hand there is nothing to prevent me from giving a brief definition of either. It does not have to be a complete definition: I may say that entropy is $\int dQ/T$ between certain limits of integration, or that a synchrocyclotron is a cyclotron in which the frequency is modulated so as to overcome the obstacle arising from the change of the mass of the nuclei with their speed. It may be objected that a person who does not know in advance what these words mean is unable to profit by the discourse. This view fails to take account of the fallibility of human memory. The listener may have forgotten what the words mean; he may even be able to recover the meanings during a few seconds of groping, but during these few seconds the speaker will go so far ahead that the gap cannot be closed. I have often observed that the place at which I lost contact with a speaker was the place at which he used a word which made me stop and ponder. It seems worth-while to avoid such dangers as far as possible.
    There is a sense in which physics is afflicted by what may be called jargons, though I should prefer to call them private languages. This is a phenomenon of recent years. Formerly physicists were few and far between, and one who did not make himself understood to his fellow-physicists a thousand miles away did not make himself understood to anybody. Nowadays many physicists do team work in large groups. In every such group a private language arises, characterized first of all by omissions. Relevant facts and even essential steps in an argument can safely be omitted within the group, because everybody knows them. In addition, the group invents all sorts of abbreviations, nicknames, and pet names for such things as parts of an apparatus, cosmic-ray tracks of various aspects, irregularities in crystal lattices, phenomena of hole-conduction, and even basic concepts of physics. No dictionary contains these terms; they travel by word of mouth, and often they do not travel fast enough. When they are spilled out before a meeting of the Society, disaster may ensue if they are not defined. Facility of travel and interchange of personnel are doing much to retard the development of a Berkeley language, an Oak Ridge language, a Murray Hill language, and the like; but the danger is always with us.
  10. 10.
    Style. The concept of style being vague and the teaching of style lying in the province of another profession, I confine myself to two remarks.
    Textbooks of style advise the writer, and therefore inferentially the speaker, to strive for a proper proportioning of long words with short, and (what often comes to the same thing) of words of Greek, Latin, or French origin with words of Saxon origin. Now, a scientific article is perforce overloaded with words which are both long and of Greek or Latin origin. This suggests that whenever the speaker has an option, he should choose the short word over the long and the Saxon word over the Greco-Latin. If a sentence contains such words as ferromagnetism or quantization or electrodynamics—not to speak of the atrocious phenomenological—it is really amazing how much the sentence will gain in grace and fluency if all the other words are colloquial and short. This policy also tends to bring out the necessary long word in bold relief.
    It is said that the style of our forerunners was largely formed by the King James Bible, and that the style of our contemporaries is influenced by The New Yorker. Neither of these publications can have much influence on those who do not read them. The suggestion is that physicists should not confine their reading to their professional literature. Read novels; read poetry; read essays; read history as written by notable writers; read Winston Churchill; and read Rebecca West—or if you simply will not go beyond the writings of scientists, read the Braggs and Eddington and Jeans and Bertrand Russell. Failure to observe this precept is partly accountable for the fact that it is seldom possible to tell from the style of an article in The Physical Review who wrote the article, and for the further fact that scientists who try to write something for the general public so often do it badly.
  11. 11.
    A suggested experiment. I have proposed, inter alia, that a speaker should speak slowly, show his slides slowly, define his private-language terms, and repeat his main points. To anyone who deprecates this advice I suggest the following experiment.
    Choose an article in The Physical Review; let it be in your own field if you will, lest the result of the experiment be too frightful. Sit down in an uncomfortable chair, and read the article—but read it according to the following prescriptions. Read straight through from beginning to end at the rate of 160 to 180 words per minute. Never stop to think over anything, not even for five seconds. Never turn back, not even to refresh your memory as to the meaning of a symbol or the form of an equation. Never look at an illustration until you get to the place where it is mentioned in the context; and when you get to that place, look at the illustration for ten or fifteen seconds and never look at it again. If this is not the way that your listeners will apprehend you when you give a paper, you are an outstanding speaker.
3.4 How Not to Listen to a Speaker
Author: William B. Bean
Original Source: The Journal of Irreproducible Results, Vol. 7, No. 2 (1959).
No orator, no matter how much energy he possesses, has a chance of defeating the drowsiness of the audience. Everyone knows that sleep during a long speech is significantly deeper than the state of hypnotic numbness known as "dozing." After such a sleep, you wake up refreshed. You have rested well. You firmly know that the evening was not wasted. Few of us have the courage to sleep openly and honestly during an official speech. After a thorough investigation of this question, I can present for the reader’s consideration several original methods which have not been published until now.
Sit in your chair as deep as possible, tilt your head slightly forward (this frees the tongue so it hangs loosely without obstructing breathing). Loud snoring drives even the meekest speaker out of his mind, so the main thing is—avoid snoring; all breathing pathways must be free. It is difficult to give precise instructions on maintaining balance in sleep, but to prevent your head from wobbling from side to side, prop it up with your two hands and your torso, forming a sturdy tripod—Archimedes knew this was a very stable device. This reduces the risk of falling onto the floor (and climbing out from under the table usually happens amidst very unpleasant animation from the public). This way, your head won’t drop onto your chest, and your jaw won’t drop. Closed eyes should be hidden in your palms; at the same time, your fingers should squeeze your forehead into a frown. This produces the impression of intense mental work and somewhat puzzles the speaker. Screaming during nightmares is possible, but this is a risk one must take. Wake up slowly, look around, and do not start applauding immediately. This might turn out to be inappropriate. Better yet, wait until they wake you up with the final applause.
3.5 Reports I Have Read… And Perhaps Written
Author: Dwight E. Gray
Original Source: Physics Today, Vol. 13, No. 11 (November 1960).
The technical report as a specialized form of scientific literature has risen from obscurity in recent years to occupy an exceptionally important position in the field of information exchange. Ideally, the main task of every report is to transmit scientific information—to transmit it accurately, neatly, and unambiguously. Alas, in practice, all "reports" break down into a multitude of categories based on their degree of approximation to this ideal, as well as to specific reasons that prevent them from achieving it. Below are described and classified the most important categories with which the author has had to deal during his fifteen years of activity in the field of scientific information. Regarding the most senseless reports, it should be said that they do not break down into clear, mutually exclusive groups. I chose the names based on the dominating feature, but on the whole, such reports are "not a proud people" and easily adopt bad habits from one another.
1. The Mystery Report
When reading such a work, it seems that the author almost intentionally tries to keep the reader in ignorance regarding what, exactly, it is written about as long as possible. In some cases, the success is complete, and the riddle remains unsolved to the very end. The mystery, of course, thickens even more if there is an absolutely uninformative title on the document. I saw two such titles recently while leafing through an abstract journal: "Review of Work" and "Current Work. Report XXIX".
Mystery reports usually begin with some historical statement like: "In 1927, Professor K. K. McJillicuddy, working at ABC University, discovered such-and-such." This initial phrase sounds approximately the same as "In a faraway kingdom, in a certain state…", with which all children’s fairy tales begin, and it requires about the same mental effort from its inventor. A similar beginning is, of course, an excellent exit for the author of the report if he is not very sure about what he wants to say, or if he is simply too lazy to start dirtying paper. For the reader, such a beginning introduces an element of mystery from the very first sentence. Beyond the initial assertion, further digressions on a historical theme may follow, and behind them other materials, only indirectly related to the true purpose of this document, which hangs in the air. And the reader plunges into the labyrinth of the report, having not the slightest idea of whether the arrows leading to the exit door actually exist, and not being even sure that this door really exists. As a rule, the reader eventually digs down to what the author was trying to tell him. However, this feat usually requires such a tension of analytical abilities and such mental subtlety that it no longer delivers joy to the completely exhausted reader.
2. The "Say It Again, Please" Report
At the disposal of anyone who wishes to learn how to write reports of this type, there is a wide selection of effective methods. First of all, the author, of course, must beware of simple non-propagated sentences. The unholy trio of subject, predicate, and object is too direct a means of transmitting information, and therefore does not satisfy. Instead of this, the author everywhere, where only possible, must use long, ornate, flowery phrases, diluting them with introductory words and starting with subordinate clauses. I will allow myself to cite as an illustration a short excerpt from a dialogue that took place several years ago at a meeting of one of the Congressional committees. An official of the Ministry of Defense was asked if their department plans the construction of a certain underground structure. Here is what he answered:
"We are trying to maintain a balance between stationary, as they are sometimes called, installations, the creation of which in individual cases may be associated with the construction of underground structures, on the one hand, and the effectiveness of our defensive means, on the other hand, which, obviously, according to the views adopted at the present time, is considered as the strongest argument against the construction of closed defensive installations."
After which one of the committee members stated: "You are simply magnificent. I didn’t understand a damn thing in this verbal eruption. For God’s sake, what did you mean?"
This example is good not only because it contains a brilliant specimen of the style "Repeat, please, one more time." The reaction of the "consumer" to such semantic twists is also completely typical.
Not wishing to create the impression that only government officials possess a monopoly on such sayings, I will tell about a less official case. The quotation I cite below is borrowed from the article "Ascending and Descending Air Currents."
(First sentence) "While the speed of the flow remains at an average level—lower than in the adiabatic case for dry air, and higher than in the adiabatic case for air saturated with water vapor, we can easily imagine that an isolated air mass, which is saturated to absolute humidity at a temperature slightly exceeding the temperature of the surrounding air, will turn out to be in a state to begin an ascending movement, since under the postulated conditions its temperature at any achieved height will be higher than the temperature of the surrounding air."
(Second sentence) "However, the opposite process—a descending flow—is not so easy to understand."
To the credit of the famous encyclopedia from which this quote is taken, it should be noted that in the next, revised edition, the second sentence was redone.
To achieve the worst results, the author must strive to maintain two numerical coefficients at an extremely high level—the "number of syllables per word" and the "average number of words per sentence." In the first case, follow the motto: "Never use a monosyllabic word if there is a synonym of six or seven syllables." Your report will especially win if, in addition, you use these long words out of place.
There are many ways to maintain the second index at a high level. For example, "bearing in mind the fact that" is always better than simply "because," "equal in order of magnitude" is much more effective than "about." You are free to start at least every sentence with such common (and such meaningless) turns of phrase as "it should be noted that," "it is not uninteresting to pay attention to the fact that," etc. These verbiages give everything stated such an impersonal character, as if removing any responsibility from the author for their content. A person, conscientiously following these recipes and supplementing them with the ubiquitous use of the passive voice and professional jargon everywhere, where only possible, and flavoring the whole dish with non-germane definitions and additions, can be completely calm about the result.
3. The "Missing Thought" Report
When reading such reports, tests on ingenuity immediately come to mind, with which we were tormented in school years. Only instead of missing words, the reader must restore missing thoughts, that is, significant chunks of information and argumentation which the author did not consider necessary to include in the text. In such cases, the reader, trying to independently make that step which appeared obvious to the author, most often finds himself in the position of students at a lecture on mathematics, about which I recently read. The professor, standing at the blackboard, was immersed in a lengthy derivation. At some point, he pronounced the standard phrase "from here it clearly follows the following" and wrote a long and complex expression, absolutely unlike anything written earlier. Then he hesitated, a puzzled expression appeared on his face, he muttered something and went out of the auditorium into his office. Appearing from there half an hour later, he explained to the audience with a satisfied look: "I was right. It is, indeed, completely obvious."
Reading such a report, you, figuratively speaking, slide smoothly along a smooth rail track following the train of the author’s thought and suddenly stumble upon a destroyed section or a precipice. You clearly see that the track continues on the other side of the abyss, but there is no bridge, and there is nothing to build it from.
4. Camouflaging Reports
Characteristic features (they can occur all at once or in certain combinations):
  1. 1.
    The results are not presented fully.
  2. 2.
    The conclusions obtained do not follow from the results.
  3. 3.
    The recommendations cited do not follow from the conclusions.
In finished, unmistakable samples of camouflaging reports, the author, pretending to be a scientist, acts in reality like a hawker. He wants to sell something in the literal sense or seduce the buyer with some idea of his. Camouflaging reports of a high class can be recognized immediately by their beautiful binding, often even with gold lettering, by the magnificent illustrations not less than in four colors, interleaved with tissue paper, and by the glossy expensive paper.
On this, perhaps, one can finish the discussion of typical "freaks" in the family of technical reports, the main task of which, as I said in the introduction and repeat now, is the transmission of information—transmission that is accurate, neat, and unambiguous.
Methodology and Education
4.1 How to Conduct Scientific Research
Authors: A. B. Mishaym, S. D. Adam, E. F. Onyakh, J. G. Bamada
Original Source: The Journal of Irreproducible Results, Vol. 6, No. 1 (1958).
Pre-publication
We propose a very useful method that allows one to publish more frequently. One must anticipate the results of an experiment and publish them in advance. This saves a tremendous amount of time. In this way, one can even save oneself the trouble of finishing the experiment; since the article is published, one can move on to something else. This trick, combined with a well-developed imagination, allows one to publish a large number of experimental papers without conducting any experiments at all, thereby saving a heap of government funds. True, a slight awkwardness may arise if someone else has already conducted the experiment and obtained different results. But an experienced scientist in this case can: a) completely ignore this circumstance; b) write a series of articles devoted to describing the subtle differences in experimental conditions that led to the difference in results; c) express deep gratitude for the pointing out of the error and write a series of articles on new experiments yielding the correct results, while using the old, erroneous ones to demonstrate all the difficulties and subtleties of this beautiful work.
Joint Publication
Much research has been devoted to the art of placing one’s name at the head of the list of authors, but some subtle questions remain unilluminated. The Alphabetical Trick. Since alphabetical order in compiling lists of authors is gradually becoming generally accepted, it is useful to concentrate on creating a solid advantage for oneself personally. This can be achieved in two ways: change your surname so that the new one begins with the letter A, or choose co-authors with surnames from the lower half of the alphabet (see, for example, G. C. Wick, A. S. Wightman, E. C. Wigner, Phys. Rev., 88, 101, 1952). But in this case, it is easy to miss the mark. One should not chase co-authors of too large a caliber. An article by A. Bar-Nudnik and Albert Einstein will always be referred to as "Einstein et al.", regardless of the order of names.
Private Correspondence ("If we can’t beat them, let’s join them")
If you have learned in a roundabout way that someone is finishing excellent work and is about to publish it, send him a letter outlining his work in the form of an idea that "recently occurred to you." Explain that you are writing to him because you heard that he is also interested in this, and later, having learned that he came to the same results "independently of you," propose a joint publication.
Congresses
The art of traveling from one international congress to another, reporting everywhere interesting work performed by someone else in your institute who for one reason or another could not go, has reached a high level of development. An expert congressman can virtuously prevent the possibility of sending anyone else on a business trip, even if he himself has not worked for many years and does not understand the works being presented.
Polemics
One must learn to use the mistakes of one’s colleagues to increase the number of one’s own printed works. The exponential growth in the total number of scientific publications is accompanied by a huge increase in the amount of nonsense appearing in so-called serious scientific journals. Without any difficulty, one can find in the literature an article that is either a) completely erroneous, or b) contains correct results obtained by committing an even number of mutually canceling errors, or c) is full of minor inaccuracies. It can be used in one of the following ways:
  1. 1.
    Write several short notes to different journals pointing out errors and inaccuracies.
  2. 2.
    Write a long article criticizing the original work and redoing everything "properly." The true difference may lie in the removal of a few insignificant sloppinesses.
  3. 3.
    Correct and rewrite the original article and publish it, citing the first as an independent but weak attempt undertaken with unsuitable means.
4.2 The Computer Poet
Original Source: Folklore / Press Releases regarding the RCA-301 computer (circa early 1960s).
The RCA-301 computing machine was taught to write blank verse. The vocabulary of the semiconductor poet is 130 words. The size of the verses is rigidly set. Starting the next poem, instead of a title, the machine puts a serial number: "Poem Number So-and-So," and at the end puts its signature: "RCA-301". Below is a literal translation of one such poem.
Poem No. 929
While blind sleep swam over broken hopes,
Space oozed with pain over broken love,
Your light was slowly banished from secretive people,
And the sky did not sleep.
RCA-301
According to programmers, this work strongly resembles the poems of modern poets Eliot and Cummings, but none of them can compete with the machine in productivity. The RCA-301 writes 150 quatrains per minute.
4.3 Preparation of Coffee in Research Institutions
Author: A. Kohn
Original Source: The Journal of Irreproducible Results, Vol. 8, No. 1 (1959).
A significant part of the time in laboratories and offices is devoted by scientists to the preparation and consumption of coffee (in Great Britain they prefer tea). The amount of beverage consumed and the time spent on consumption range from one cup and five minutes a day to several dozen cups and several hours. There is a very acute need to regulate this activity. This note represents the first attempt to generalize the rich experience currently going to waste.
Materials and Methods
All commercially available varieties of coffee are used (with the exception of acorn and synthetic), of fine, medium, normal, and coarse grind, as well as in beans. Boiling is performed in vessels made of glass, aluminum, or stainless steel, including laboratory beakers, fireproof flasks, strainers, autoclaves, and (there was once a case) distillation stills. However, ordinary coffee pots are also used. Sources of energy for raising the temperature of the extracting water can be the flame of a gas burner, electric current, superheated steam, exhaust gases of an internal combustion engine, and the reaction of alcohol oxidation (do not use ether – it is explosive). In summer, cooling systems are used: refrigerators, freezers, ice cubes, dry ice, and liquid air.
Methods
  1. 1.
    "Amateur". Coarse-ground coffee is poured into cold raw water, then the water is brought to a boil. The sediment is thoroughly shaken, and the resulting suspension is poured into cups. That the resulting product is indeed coffee can perhaps only be verified with the help of photometric measurements. The same applies to a slightly more economical modification of the amateur method, where the sediment is filtered off and saved for reuse.
  2. 2.
    "Professional". Water is heated to $99^{\circ}$C, fine-ground coffee is added (approximately 1.2 grams per cup), the liquid is brought to a boil and removed from the fire. Centrifuging the sediment is not necessary. The resulting drink has a delicate taste, which sometimes can be killed by adding milk.
  3. 3.
    "Expert". Statistical processing of tasting results allowed for an important improvement: water is first brought to a boil, then the flame is reduced to a minimum, and coffee is poured into the water (one spoon per cup). Everything is boiled for 10 seconds, then settles for a few minutes, and one can drink.
  4. 4.
    "Express" (should be distinguished from the "Expresso" method described below). Instant coffee is used in combination with boiling or hot water according to the instructions written on the tin. This method has two advantages over all others: speed and the absence in this case of the notorious "wonderful aroma of fresh coffee," which allows avoiding an invasion of the thirsty from the neighboring laboratory.
  5. 5.
    "Expresso". This name has already become a household word for steam extraction, when superheated steam is passed through compressed coffee powder and then cooled. The resulting condensate possesses the color and smell of coffee. In laboratory conditions, one can use the "Soxhlet" extraction apparatus, however, this method is too labor-intensive for use, except in cases where you have a surplus of technical personnel, a lack of ideas, and you cannot think of a more reasonable way to use the apparatus.
4.4 Advice to University Professors
Author: Marvin Camras
Original Source: IRE Transactions on Audio.
In our days, educators are constantly revising curricula, and we often hear about "new math" and other innovations. We propose here to introduce small lecture courses into the curricula, which, from our point of view, would be extremely useful for future engineers.
Anti-standardization (Creative Invention)
The goal of the course is to teach the creation of devices in which not a single part can be replaced by a standard one. This requires great ingenuity, however, successful labor is generously rewarded. The expensive and wasteful "Anti-standard" is a supreme achievement that is rarely obtained by accident. A scientific approach in this matter allows creating super-non-standard devices in which all dimensions are non-typical, all parts are electrically, mechanically, and chemically incompatible, possess increased corrosion instability and increased brittleness, and thus possess the maximum rate of failure.
Committees
The goal of the course is to teach the future engineer the technique of using committees and working in them. As is known, committees are the ideal means for eliminating oneself from any responsibility, for delaying the execution of all tasks, and for creating a "managerial" mood among directors. Additionally, a committee meeting is a not-bad means to kill evening time intended for rest. This cycle of lectures will help listeners hone their art of putting off things from day to day, appearing smarter than they are, teaching them to use jargon, and in the role of chairman will help to amaze everyone with their concentration, joke effectively, and gracefully close meetings.
Ability to Adapt to the Environment
Previously, everyone relied exclusively on intuition in this area, and it only recently became a science thanks to a number of selfless enthusiasts. Recognized masters of this business should be invited to give the course. They will share experience in their art of appearing eternally busy, confident in their abilities, full of infectious enthusiasm, striving to expand this or that department 1) by personnel, 2) territorially, 3) in the sense of financial appropriations. They will teach how to effectively plan a budget, how to report on expenses with brilliance, how to process bills, appear smart in the presence of 1) engineers, 2) administrators, 3) cleaners, and also the ability to appear unambitious. They will give advice on how to eat, what to drink, how to choose a car, a wife, and a lawnmower.
Information Channels
The goal of the course is to teach the future engineer methods of approaching the boss’s secretary, viewed as the most important source of useful information. However, without theoretical preparation, one can do foolish things, failing to take into account certain subtleties, like the fact that the secretary from another department, with whom the boss’s secretary drinks her coffee during the break, is possibly more talkative.
4.5 Advice to the Examiner
Original Source: Electronics.
  1. 1.
    First of all, explain to the examinee that his entire professional career may collapse because of his unsuccessful answer. Emphasize the importance of the situation to him. Put him in his place from the very beginning.
  2. 2.
    Ask the hardest questions right away. If the first question is sufficiently difficult or confused, the examinee will get too nervous to answer the following questions, no matter how simple they are.
  3. 3.
    Addressing the examinee, maintain restraint and dryness; with other examiners, however, be very cheerful. It is effective to address other examiners from time to time with mocking remarks regarding the examinee’s answers, ignoring him himself, as if he is not in the room.
  4. 4.
    Force the examinee to solve problems by your method, especially if this method is unusual. Limit the examinee by inserting a multitude of instructions and reservations into every question. The idea here consists of complicating a task which without this would be quite simple.
  5. 5.
    Force the examinee to make a trivial error, and let him rack his brains over it as long as possible. Immediately after he notices the mistake, but just before he understands how to fix it, contemptuously correct him yourself. This requires high insight and precise timing, which is achieved only by great practice.
  6. 6.
    When the examinee begins to drown, never help him scramble out. Yawn… and move on to the next question.
  7. 7.
    Ask the examinee questions from time to time like: "Didn’t you cover this in elementary school?"
  8. 8.
    Do not allow the examinee to ask clarifying questions and never repeat your own explanations and statements.
  9. 9.
    Every few minutes ask if he is worried.
  10. 10.
    Wear dark glasses. Impenetrability makes people nervous.
  11. 11.
    Finishing the exam, say to the examinee: "Wait outside the door. We will call you."
4.6 Mathematization
Author: N. Vanserg (Pseudonym of Harvard Geology Professor H. E. McKinstry)
Original Source: The American Scientist, Vol. 46, No. 3 (1958).
In an article published several years ago, the author already hinted (with all the subtlety befitting the case) that, since most scientific truths (if one sorts them out) are relatively simple, any self-respecting scientist must, for the purpose of self-defense, try to prevent his colleagues from understanding that his own ideas are also simple. Therefore, if you manage to give your publications a sufficiently incomprehensible and uninteresting form, no one will attempt to read them, but everyone will bow before your erudition.
What is Mathematization?
A further development of the currently well-known methods of writing articles in a language only approximately resembling English can serve the subtle art of using mathematical symbols wherever possible. This art has only one drawback: what if a clever rogue is found, no worse than you skilled in these primitive tricks, who figures out the confused argumentation and discovers the hidden simplicity? Fortunately, there are many ways to nip such a vile attempt in the bud.
Typographical Tricks
The most ancient method is to write the wrong letters in formulas, for example $\phi$ instead of $\psi$. Even the simple placement of the symbol $\exp$ to the right of the brackets is sometimes capable of working wonders. This trick is conscious fraud, but it rarely entails punishment, since one can always shift the blame to the typesetter. Generally, the author, as a rule, does not have to work on inventing such traps, for typists and formula-inserters finely feel the needs of authors and themselves show initiative and good will in this regard. One can fully rely on them and then just not read the proofs.
Strategy of Secret Symbols
But what if, thanks to an accidental concurrence of circumstances, the formulas escaped distortion beyond recognition? After all, the reader, if he knows what each symbol means, will figure them out. This is where the front line of your defense passes: make it so that he never learns this. For example, you can print in black and white in a footnote on page 35 that $V_{1}$ is the total volume of phase 1, and on page 873 calmly introduce it into an equation. And your conscience will be absolutely clear—after all, you did say eventually what this symbol means. By secretly introducing all letters of the Latin, Greek, and Gothic alphabets, you can force the curious reader, interested in some paragraph, to read the entire book in reverse order to find out the meaning of the notations. Books that read from end to beginning no worse than from beginning to end produce the greatest impression.
When reading backwards becomes a habit for the reader and he considers this method normal, confuse the tracks. Insert, for example, $\mu$ into an equality on page 66, and wait with its definition until page 86…
But the moment has come when the reader thinks he already knows all the letters. It is high time to use this fact to besiege him a little. Every schoolboy knows what $\pi$ is, and this will help you break away from the enemy again. The poor reader will long automatically multiply everything by 3.1416 before he understands that $\pi$ is osmotic pressure. If you are careful and do not blab ahead of time, this will cost him about an hour and a half. The same principle can, of course, be applied to any letter. Thus, you can on page 141 absolutely honestly write that $F$ is free energy, and if the astute reader is used to the fact that $F$ is free energy in Helmholtz’s definition, then he will spend a mass of his own free energy deciphering your equations before he understands that you meant Gibbs free energy all along, which the reader thinks is $G$. generally, $F$ is a wonderful letter, it can denote not only any free energy, but also fluorine, force, farad, as well as a function of an arbitrary number of real and complex variables, thereby significantly increasing the degree of randomization dSdS ($S$, as is known, denotes entropy and… sulfur).
Asterisks and numerical indices, which denote footnotes at the bottom of the page, can also be used for military ruses. Designate, without saying a bad word, some pressure as $P^{*}$, so that the unsuspecting reader searches for a footnote, which on this page, needless to say, is not there at all. And when the seeker of truth reads that $S$ constitutes $10^{14}$ cal, he will think: "Wow, what a hell of an abyss of calories!"—and will continue to think so until he reads the page to the end, stumbles there upon footnote number 14, and says: "A-a-h…".
"Therefore"
But the greatest success is achieved with the help of the following trick: from the finished manuscript, you tear out two pages of calculations, and in their place insert the word "therefore" and a colon. I guarantee that the reader will guess for a good two days where this "consequence" came from. Even better is to write "obviously" instead of "therefore," since there is no reader who would dare to ask anyone for an explanation of an obvious thing. By this, you not only confuse the reader but also instill in him an inferiority complex, and this is one of the main goals.
All that has been said is, of course, elementary and generally known. The author is currently finishing a two-volume work on mathematization, including examples and problems for independent exercises. There will be so many classified notations, riddles, typos, and consequences flowing from nowhere in it that no one will be able to read this work.
4.7 Instructions for Authors
Author: Jack Ewing
Original Source: The Journal of Irreproducible Results, Vol. 12 (1963).
In the journal that I head, as a rule, articles are accepted that cannot be pushed through anywhere else. If your article was returned from another journal (with a small letter), iron it with an iron and send it to our Journal (with a capital letter).
Formatting the Manuscript
Text: The manuscript must be typed on standard sheets of parchment size $81/2\times 11$ inches. Typing should be done on no more than two sides of the sheet, starting every paragraph from a new page. Leave margins of $4\dots 41/2$ inches on both sides of the page.
The author must use clear, transparent, crystal English language, preferably using not less than two-syllable words consisting of not more than four letters. One should not use words understandable to your colleagues. For example, do not write "shortened", but write "reduced", not "changed", but "modified". Do not use sentences longer than 120 words without including in them at least one verb or participle.
In everything concerning spelling, use of capital letters, and other things, follow Webster’s dictionary. The editor will redo everything anyway. Abbr. s. b. red. to min. Do not make tables from data that can be listed in the text. Do not list in the text data from which a table can be made. Do not express ratios in mg/kg.
References: Name the author, his address, and the volume number. If possible, the year. (Instead of authors’ surnames, their nicknames can be cited if they are generally known.)
Illustrations: They can be made in various ways. Special clarity is not required. On the back of each photograph, briefly outline instructions for the editor (avoid unprintable words).
Conclusions: If possible, conclusions should be shorter than the main text and presented in a form allowing their use as abstracts for abstract collections.
Not Only Physicists Joke
5.1 Parkinson’s Law in Scientific Research
Author: C. Northcote Parkinson
Original Source: New Scientist (Adapted).
To an outside observer, it seems that people who apply to Funds and Trusts, to organizations bearing the names of Rockefeller, Guggenheim, and Ford, before doing so, should investigate the question of chasing subsidies from a scientific point of view. In the contrary case, disappointment awaits them. Knowing that savings are necessary, and expenses must be cut, they often think that it is only worth coming up with a plan, drawing up a budget, and you will be received with open arms.
Let us assume that a certain Dr. Grant has a plan: to determine the incidence of philately in Hong Kong. In his dreams, he sees himself in the office of the Vanderfeller Trust, before the Board of Directors, which includes Dr. Please, Mr. Giveall, Mr. Distributeall, and Mr. Riskall. They are delighted with his plan and are worried only about whether the doctor has enough with those one and a half million dollars he is asking for. They develop the impression that five million is a more real figure. "What dollars do you have in mind?" asks the timid Dr. Grant. "I meant, of course, American dollars," answers Dr. Please. Mr. Giveall writes out a check and wishes Dr. Grant all the best. The conversation is over. Such are dreams.
But what happens in reality? Dr. Grant finds himself in an office face to face with Dr. Stingy, Mr. Zilch (Mr. Fig-vam), and Mr. G. de Garantia. Stingy says that philately is a social evil rather than a disease, and therefore goes beyond their competence. de Garantia considers the entire plan politically dangerous, since it concerns Hong Kong. All together explain that this whole idea is inadmissible, unacceptable, immoral, and illegal. Dr. Grant is thrown out into the street, and the doorman is instructed not to let him in henceforth. The papers left by him are sent to the prosecutor as material evidence for initiating a criminal case on charges of extortion.
You ask, what is Dr. Grant’s mistake? There are people who want to spend money. And he offers a plan on how to spend it. So why not? His project, rejected so categorically, is no more useless than many others already implemented. Why was he kicked out? Simply because that is his plan. That is exactly exactly why they recognized it as brilliant. The main thing in the art of obtaining subsidies and grants is to convince the officials of the Fund that the project was proposed by them, and you are merely an obedient pawn, agreeing in advance with all their suggestions.
Let us assume now that you received the subsidy—maybe from the government, maybe from public charitable organizations, or most likely from a private donation. Now your immediate task is to spend and even overspend these funds as quickly as possible, in order to ask for more next time. Benefactors prefer expenses on the construction of a building—after all, a memorial plaque with the family name of the donor can be solemnly laid in its foundation, and hung on the wall. And what advertising can Dr. Grant provide them? A tombstone on his own grave?
If you decide to erect a building, it is best to attach a pompous memorial arch to an already existing hospital between the laboratory and administrative buildings, and to build a cozy apartment for yourself somewhere nearby: it is very important to live close to the place of work. Furthermore, the inscription on the memorial plaque can be composed in such a way that it creates the impression (without asserting it directly) that the donor paid for the entire hospital. The composition of the inscription should be entrusted to a specialist in ambiguities. For a small, or rather relatively small fee, I myself would undertake to compose such an inscription.
Building a grand entrance is the surest way to attract donations (and remember—a hospital can have more than one gate, but several). But all buildings have a common drawback: the number of scientists working in them increases very rapidly, they fill the building and overflow it, as a result of which the problem of space arises more acutely than before. It seems that the premises themselves shrink in size (this, of course, is an optical illusion, but the fact noted above is not subject to any doubt).
A similar law of continuous growth applies to the number of journals covering progress in a certain field of science. Why does this happen? For a long time, no one could give an intelligible answer to this question, so I am not without pleasure to report that I have managed to uncover the true causes of the multiplication of scientific journals and laws by which it occurs.
I will explain these laws with the following example. Let us suppose that the oldest and most respectable of all journals on clinical medicine (Journal No. 1) was published for many years by Professor A. This professor was an outstanding person (so outstanding that many probably guess his name, which I do not dare to name here). He died several years ago. If A was mistaken in anything (and who among us does not make mistakes!), it was only in that he refused to publish all articles with which he did not agree. In practice, this means—all articles written slightly above the student level. This lasted for several years and terribly bored Professor B, who never agreed with Professor A.
If, for example, they were asked to write the word "vinaigrette", I am sure that they would write it differently. With such a striking difference in views, it is not surprising that the articles of Professor B for twenty-three years were invariably returned to the author. Upon the expiration of this term, he decided to found Journal No. 2. This publication began to come out on a more liberal basis, and at first, everything was printed in it, except for the works of those authors regarding whom it was accurately known that they are followers of Professor A. But Professor B had his own high principles. He believed that any views, including those which differ slightly from his own, deserve the right to free expression; he insisted only that they be set out consistently and scientifically. And so, one day, he had to reject the works presented by Professor C. (I must speak with caution: he is alive and well and deservedly receives a pension.) Everyone considered him an original and interesting thinker, but found that he was somewhat hasty in his conclusions and slightly careless in the presentation of results.
Having discovered that his articles are rejected by Journals No. 1 and No. 2, he became the founder and first publisher of Journal No. 3, which did not refuse the most unintelligible works on the vaguest topics. Everyone knows what journal I have in mind. But if you know, then you must notice that he also has a reputation that he values. His literary level is very high. Perhaps nothing is communicated in his reports, but the drawings prove the assertions that they must illustrate, but the grammar in this journal is above all criticism. Following his clinical advice, you can become a murderer, but the pages of this publication were never defiled by clumsy phrasing. Feeling obliged to protect the literary reputation of the journal (only that!), the editor was forced to reject the works of Professor D.
But we know all this—Professor D is not such a person who can be shut out of the printed pages. And here the reader receives Journal No. 4. But even D must draw a dividing line somewhere! He stubbornly refuses to publish the works of Professor E on the pretext that E does not know spelling. (And this, frankly speaking, is true.) Of course, some will argue that the article can be finalized in the editorial office. (And this, of course, is fair.) But Professor D can also be understood, and I would not blame him for narrow-mindedness. He simply does not want it to be rumored about Journal No. 4 that they accept everything that is typed on a typewriter on one side of the sheet at double intervals. He must maintain the prestige of the journal. On the other hand, no one will raise a hand to throw a stone at Professor E for the fact that he started publishing Journal No. 5. It is precisely such a development of events that led to the fact that only on questions of dentistry and prosthetics, we publish about eighty journals.
If progress in science were measured only by the quantity of published works, then the number of existing journals could become a source of satisfaction and pride. But it is necessary to remember that every journal needs an editorial board and an editorial staff, several editors with assistants, numerous reviewers, consultants, and critics. Due to the man-hours spent on academic journalism, a mass of time is lost, intended for scientific work. If everyone involved in one way or another in a certain issue read the journals published by others (and this is the best way to avoid duplication), then it is clear that they would not have time for anything else. Interestingly, those few people whose research represents at least some value usually keep each other informed of their affairs through personal correspondence.
From what has been said, one can erroneously conclude that any person who has devoted himself to scientific work ends his career as an editor. This is incorrect. Only those who fail to take an administrative position become editors.
What is the normal course of events? A person who has made a significant contribution to science is persistently offered subsidies to expand the front of research. This is exactly what happened with Dr. Hodgkins, a brilliant employee of Professor Wilkins. Can one forget his speech delivered at the meeting of the American Federation of Clinical Research in 1938! According to his theory, artists creating modern abstract paintings, as a rule, suffer from color blindness, and in some cases—dementia. With this, he created a reputation for himself, and the Dalvzylkins Foundation hurried to generously subsidize his further work. Hodgkins was asked to find out whether composers writing dance music for young people really lack a musical ear (as Professor Wilkins suspected) or whether they are simply mentally underdeveloped (the opinion to which Hodgkins himself leaned). This was a grandiose project. Sector A was intended for work with artists suffering from color blindness, and Sector B—for examining mentally deficient jazz players.
From now on, Dr. Hodgkins had to deal with the organization of the work of his staff, numbering 432 people, of whom 138 had medical or scientific qualifications, 214 had secondary and higher technical education, and 80 were hired for clerical work. Well, and Dr. Hodgkins himself lost the opportunity to engage in scientific work—obviously. But not many people understand that on this path they also lose the opportunity to manage someone else’s scientific work. They will spend all their time on problems of rational use of working premises, dealing with safety engineering, compiling vacation schedules, streamlining wages, etc., etc.
Now we can formulate the **"Parkinson’s Law for Scientific Research"**. Here it is:
"Success in scientific research causes such an increase in subsidies that the continuation of research becomes impossible."
5.2 Common Sense and the Universe
Author: Stephen Leacock
Original Source: New Scientist, Vol. 13, No. 271 (1962).
"In the space of one hundred and seventy-six years the Lower Mississippi has shortened itself two hundred and forty-two miles. That is an average of a trifle over one mile and a third per year. Therefore, any calm person, who is not blind or idiotic, can see that in the Old Silurian Period, just a million years ago next November, the Lower Mississippi was upwards of one million three hundred thousand miles long, and stuck out over the Gulf of Mexico like a fishing-rod. And by the same token any person can see that seven hundred and forty-two years from now the Lower Mississippi will be only a mile and three-quarters long, and Cairo and New Orleans will have joined their streets together, and be plodding comfortably along under a single mayor and a mutual board of aldermen. There is something fascinating about science. One gets such wholesale returns of conjecture out of such a trifling investment of fact."
Mark Twain, "Life on the Mississippi"
Speaking in December 1941 at the annual meeting of the American Association for the Advancement of Science, Professor Edwin Hubble of the Mount Wilson Observatory announced that the Universe is not expanding. This was truly good news, if not for the general public, who had no reason to suspect that the Universe was expanding in the first place, then at least for those of us who humbly try to "keep up with the progress of science."
For the last twenty-five years, specifically since the day Professor de Sitter published his terrible hypothesis in 1917, we, as best we could, have tried to live in this expanding Universe, where every part flies away from every other part at a furious speed. It reminded us of that desperate lover who jumped on his horse and rode madly in all directions. The idea was majestic, but it created a certain feeling of discomfort.
Nevertheless, we had to believe it. We had to rely, for example, on the authority of Spencer Jones of the Royal Astronomical Society, who in his 1940 book Life on Other Worlds claimed that "a distant galaxy in the constellation Boötes is receding from us at a speed of 24,300 miles per second." From this, it follows that it is at a distance of 230,000,000 light-years. I will remind my friends—lovers of science—that a light-year is the distance light travels in a year, moving at a speed of 186,000 miles per second. In other words, this "distant galaxy" is located away from us at a distance of 1,049,970,980,000,000,000,000 miles…
And now it turns out that it is not moving away at all! The astronomers did not simply assume that the Universe is expanding, but proved it by studying the behavior of the red part of the spectrum, which turned even redder from such a discovery, like the shy water in Cana of Galilee that "saw the Lord God and blushed." One of the most outstanding of our astronomers, Sir Arthur Eddington, wrote a book The Expanding Universe to bring this fact to general attention. Astronomers took the news of this universal explosion with the same calmness with which they once accepted the news of the impending heat death of the Universe; according to the second law of thermodynamics, it must eventually die of cold.
But the joy delivered to us by Professor Hubble is tempered by some doubts and reflections. Do not think that I express disbelief in science or disrespect for it (in our days this would be as monstrous as not believing in the Holy Trinity in the time of Isaac Newton). But still… Today we expand, tomorrow—we shrink: first we suffer in a curved and closed space, then this loop is loosened and dissolved completely; only just were we sentenced to agonizing death at a temperature of minus 273 degrees Celsius, in a cold that grips everything and everyone,—and suddenly it got warmer again.
So do we have the right to ask—what is the matter? Where are we? Einstein answers this question: "Nowhere, because there are no places where we could be at all." So pick up your books, follow the development of science, and wait for the next astronomical congress.
Take the history with the famous Second Law of Thermodynamics, this curse of inexorable fate, which dooms the entire Universe (or at least all life in the Universe) to death from cold. Now I recall with regret the tears I shed, heartily sympathizing with that last bunch of doomed ones who are destined to die at a temperature of 273 degrees below zero Celsius, at absolute zero, when all heat is exhausted and all molecules stop. Furnaces will not burn, matches will not light, and there will be no one to light them…
I remember when I first read about this cruel law in a "popular science" book titled Our Time is Running Out as a small boy. It was written by Richard Proctor and produced a terrifying impression. The Sun, it turns out, is cooling down and will go out completely. This was confirmed by Lord Kelvin. As a true Scotsman, he feared nothing and left the Sun and the entire Solar System only ninety million years of life.
This famous prediction was first made in 1824 by the great French physicist Sadi Carnot. He showed that all bodies in the Universe change their temperature—hot ones cool down, and cold ones heat up. Thus, they equalize their temperature. It is the same as dividing a rich inheritance equally among all poor relatives; the result will be general poverty. So the cold of world space must engulf us all eventually.
True, a ray of hope appeared when Ernest Rutherford and others discovered radioactivity. Radioactive atoms, decaying, seemed to be able to maintain fire on the Sun for quite a long time. This pleasant news meant that the Sun, on the one hand, is much younger, and on the other—much older than Lord Kelvin assumed. But all the same, this is only a delay. Eventually, scientists calculate, we will still freeze.
When in place of medieval superstitions came enlightenment, the first sciences that stood out and defined themselves were mathematics, astronomy, and physics. By the beginning of the nineteenth century, everything was put in its place; the Solar System rotated so sleepily and smoothly that Laplace managed to convince Napoleon that God, who would look after it, was generally unnecessary. Gravity worked like a clock. Chemistry, which in the time of Benjamin Franklin was merely a collection of incoherent experimental data, turned into a science after Dalton appeared and showed that any substance can be broken down into very, very small atoms, atoms combine into molecules, and everything goes according to plan. Since Faraday and Maxwell, electricity took its place in the new scientific order (it turned out that it is the same thing as magnetism).
Around 1880, it became clear that the world is perfectly explained by science. Metaphysics was still muttering something in its sleep. Theology still delivered sermons. It tried to dispute many discoveries of science, especially in geology and in the new evolutionary theory of life. But science already paid little attention to this.
Because everything was very simple. There is time and space—things too obvious to explain. There is matter, made of small solid atoms, similar to tiny grains. All this moves, obeying the law of universal gravitation. Nebulae condense into stars, stars eject planets, planets cool down, life is born on them, it develops and becomes rational, first anthropoid apes appear, then Bishop Wilberforce and, finally, Professor Huxley.
There remained a few small ambiguities, for example, the question of what Space, Matter, Time, Life, and Reason actually are. But Herbert Spencer very opportunely guessed to call them unknowable, locked them in a desk drawer, and left them there.
Everything was explained by Mechanical Determinism. There remained only this repulsive skeleton in the box of the writing desk. Yes, there was also something strange and mysterious in electricity, which was not exactly a thing, but not exactly a fiction. There was still a strange riddle about "action at a distance," and electricity only exacerbated it. How does gravity get from the Earth to the Sun? If there is nothing in space, then how does light fly to us from the Sun in eight minutes and even from Sirius—in eight years? Even the invention of "ether," a kind of universal jelly, did not relieve science of a certain unconvincingness.
And so, right on the threshold of the 20th century, the whole building began to crumble. First came the warning that not everything is smooth—the discovery of X-rays. Roentgen discovered them, and since then most physicists call them Roentgen rays. Sir William Crookes, experimenting with tubes filled with rarefied gas, opened "radiant matter" just as accidentally as Columbus opened America. The British government immediately granted Crookes a knighthood, but it was already too late. The deed was done.
Then followed the works of a whole school of radioactivity researchers. The crown of them was the work of Rutherford, who revolutionized the theory of the structure of matter. I knew Rutherford well—we were colleagues at McGill University for seven years—and I can confirm that he acted without a premeditated intention to shake the foundations of the Universe. But he did exactly that, for which he was also made a lord.
One should not confuse the works of Rutherford on nuclear physics with the theory of space and time created by Einstein. Rutherford never referred to Einstein in his life. Even when he worked at the Cavendish Laboratory and, showing black ingratitude, smashed the very atoms that made him famous, even then he needed nothing from Einstein. I once asked Rutherford (this was in 1923, Einstein’s worldwide fame was at its zenith) what he thought of the theory of relativity. "Ah, nonsense!" he answered. "We don’t need this for our work!" His biographer and admirer, Professor Eve, says that when the German physicist Wien told Rutherford that not a single Anglo-Saxon understands the theory of relativity, Rutherford replied: "Naturally, we have too much common sense."
But still, the main troubles started precisely with Einstein. In 1905, he announced that there is no absolute rest. And since then it has been gone. But only after the First World War did the reading public throw itself at Einstein, and shelves in shops began to break from books "about relativity."
Einstein knocked out space and time just as Rutherford knocked out matter. The general view of the theory of relativity on space is very simple. Einstein explained to everyone that there is no such place as "here." "But I am here," you say. "Here—is just the place where I am sitting right now." But you are moving! Earth turns, and you turn with it. Together with the Earth, you move around the Sun, and together with the Sun—following the "distant galaxy," which itself rushes at a speed of 26,000 miles per second. So what is this place—"here"? How do you mark it? It all reminds one of the story about two idiots fishing. One says to the other: "Listen, we should mark the place where we caught this healthy fish," and the other replies: "I already made a mark on the boat." Here is your "here"!
Einstein’s discovery of the curvature of physical space was greeted with an explosion of applause, such as had not been heard perhaps only since baseball. The brilliant scientist, Sir Arthur Eddington, who handles space and time as a poet (even his discussions about gravity are permeated with humor: he says that the ideal opportunity to study gravity is available to a man falling in an elevator from the twentieth floor), so, Sir Arthur Eddington applauded louder than anyone. According to him, without this curvature, it is impossible to understand anything at all. We crawl like a fly on a globe, thinking it is flat. The mysteries of gravity puzzle us (I do not mean those few lucky ones who had the chance to fall in an elevator from the twentieth floor. But revelation descended on them too late, and the revelation consists in the following: we do not fall at all, we simply curve). "Admit the curvature of space," wrote Eddington in 1927, "and the mysterious force will disappear. Einstein banished this demon."
But now, fourteen years later, it begins to seem that Einstein is little worried whether space is curved or not. To him, apparently, it’s all the same. One famous physicist, heading a faculty at one of the largest universities, recently wrote to me on this subject: "Einstein hopes that the general theory, taking into account some properties of space resembling the fact that space is usually called curved, may in the future turn out to be more fruitful than it, apparently, takes place at the present time." Spoken purely professionally. The majority says simply that Einstein waved his hand at curved space. It is the same as if Sir Isaac Newton, yawning, said: "Ah, you are talking about this apple—maybe it didn’t fall at all?"
— It is necessary to admit, gentlemen, that this discovery is somewhat staggering.
5.3 Saga of a New Hormone
Author: Norman Applezweig
Original Source: The Journal of Irreproducible Results (1959).
In recent months, the world learned about the discovery of three miracle drugs by three leading pharmaceutical firms. Upon closer examination, it turned out that all three preparations are one and the same hormone. If you are interested to know how one and the same chemical compound receives several different names, let’s trace the chain of events preceding the creation of the miracle drug.
First, it is usually discovered quite by accident by a physiologist in pursuit of two other hormones. He gives it a name reflecting its function in the organism, and predicts that the new compound may be useful in treating a rare blood disease. Having processed one ton of fresh bull glands, delivered directly from the slaughterhouse, he isolates 10 grams of pure hormone and sends them to a specialist in physical chemistry for analysis.
The physico-chemist discovers that 95% of the hormone purified by the physiologist consists of various impurities, and the remaining 5% contains at least three different compounds. From one such compound, he successfully isolates 10 milligrams of pure crystalline hormone. Based on the study of its physical properties, he predicts a possible chemical structure of the new substance and expresses the assumption that its role in the organism, most likely, does not coincide with the predictions of the physiologist. Then he gives it a new name and forwards it to an organic chemist for confirmation of his assumptions about the structure of the compound;
The organic chemist fails to confirm these assumptions and instead discovers that the new compound is just a methyl group different from a substance recently isolated from melon rind, which, however, is biologically inactive. He gives the hormone a strictly chemical name, absolutely precise, but too long and unsuitable for wide use. For brevity’s sake, a name invented by the physiologist is kept for the new substance. Eventually, the organic chemist synthesizes 10 grams of the new hormone, but tells the physiologist that he cannot give a single gram, for all these grams are absolutely necessary for him to obtain derivatives and further structural studies. Instead, he gives him 10 grams of that compound which was isolated from melon rind.
Here, a biochemist joining the search suddenly announces that he has discovered this same hormone in the urine of pregnant sows. Based on the fact that the hormone is easily split by a crystalline enzyme recently isolated from the salivary glands of the South American earthworm, the biochemist persistently asserts that the new compound is nothing other than a variation of vitamin B1G, the deficiency of which causes shifts in the acid cycle in annelids. And he changes the name.
The physiologist writes a letter to the biochemist with a request to send some South American worms.
The nutritionist finds that the new compound acts exactly the same as "Factor PFF", recently extracted from chicken manure, and therefore advises adding it to white bread for the purpose of increasing the viability of future generations. To emphasize this extremely important quality, the nutritionist invents a new name.
The physiologist asks the nutritionist for a piece of "Factor PFF". Instead, he receives a pound of raw material, from which "Factor PFF" can be made.
The pharmacologist decides to check how the new compound acts on gray rats. With confusion, he is convinced that after the first injection, the rats go completely bald. Since this does not happen with castrated rats, he comes to the conclusion that the new preparation is a synergist to the sex hormone testosterone and an antagonist to the gonadotropic factor in the pituitary gland. From here he makes a conclusion that the new agent can serve as an excellent drops for instillation into the nose. He invents a new name and sends 12 bottles of drops together with a dropper to the clinic.
The clinician receives samples of the new pharmaceutical product for testing on patients with sinus colds. Instillation into the nose helps very weakly, but he with surprise sees that three of his cold-stricken patients, who were additionally suffering from a rare blood disease, suddenly recover. And he receives a Nobel Prize.